### v3

# arXiv:2511.14790v3[physics.plasm-ph] 02 Mar 2026

Abstract
Solving the kinetic equation for ionization-recombination processes in cold plasmas for temperatures much lower than the first ionization potentials, we derive an explicit expression for the bulk viscosity. We obtain that bulk viscosity can be many order of magnitude bigger than the shear viscosity. Our result for the relaxation time reveals that the Mandelstam-Leontovich approximation for the frequency dependence of the bulk viscosity is in practice an exact result for the cold plasmas. The illustrative numerical examples correspond to the plasma cocktail of the solar chromosphere at the height of the minimal polytropic index. We systemize also explicit formulae for the thermodynamics of the binary alkali-noble cocktails. The possible application for the acoustic heating of the inner solar atmosphere up to the transition region is shortly discussed together with the evaluation to confirm the theory by laboratory plasmas.
I Introduction. Bulk viscosity in different physical conditions.
Why do champagne glasses ring hollow during a toast? What is the dominant mechanism for sound absorption in ocean water? Can we do a similar experiment in a glass of Epsom salts? Are models of solar corona heating with acoustic waves realistic? There is a huge number of physical systems in which bulk viscosity prevails over shear viscosity. But due to the complexity of the problem, we are unable to calculate bulk viscosity from first principles. Except perhaps, for the simplest case of cold plasma, which is the subject of our research. Our purpose is to derive a convenient formula for the bulk viscosity applicable for the case when temperature $T$ is much smaller than ionization potential $I_{a}$ of the main atomic ingredients of a plasma cocktail.
I.1 Motivation. Acoustic heating in the inner solar atmosphere
In some sense our study is inspired by many articles where acoustic heating of the solar atmosphere is considered [1, 2, 3, 4, 5]. That is why for an illustration we give as a numerical example corresponding to the solar chromosphere at high above photosphere $h=1894$ km from the height dependent solar atmospheric profile [6, Model C7] (AL08). When hydrogen dominates, it is convenient to parameterize the chemical compound by the abundances of other elements $\overline{a}_{a}$ given in Ref. [6, Table 2]. In such a way per every hydrogen atom we have $\mathcal{N}_{a}$ atoms
$\mathcal{N}_{a}=\sum_{a}\overline{a}_{a}\approx 1.1.$ (1)
The corresponding concentrations of chemical elements are correspondingly $\overline{a}_{a}/\mathcal{N}_{a}$. In such a way, if $n_{\rho}$ is the number of hydrogen atoms per unit volume, the number of other atoms per unit volume is
$n_{a}=n_{\rho}\,\overline{a}_{a}.$ (2)
I.2 A short discussion on the bulk viscosity
In the air we breath the energy transfer between translational and intramolecular degree of freedom is relatively slow and he first and second viscosity are comparable $\zeta/\eta\sim 1$. Perhaps Albert Einstein was the first who paid attention on the theory of bulk viscosity in partially dissociated gases [7]. The direct translation of this paper is Sound propagation in partially dissociated gases and in this sense, our study is a continuation of Einstein’s idea realized in partially ionized atomic gases with the main focus being the solar chromosphere.
Calculating the bulk viscosity from first principles has not been yet presented, as what is known to us so far. In special cases in the presence of strong magnetic fields, rotations [8] and in molecular dynamics [9, 10, 11] bulk viscosity analysis has been made, but in general the bulk viscosity is often neglected when treating plasma problems.
For the importance of the bulk viscosity $\zeta$ in the physics of plasmas we wish to point out the recent study by Istomin & Kustova [12] who studied transport coefficients and heat fluxes in non-equilibrium high-temperature flows with electronic excitation. They confirmed the intuitive conclusions by Einstein and calculated that chemical processes can lead to bulk viscosity domination $\zeta\gg\eta$. The authors of this study performed a non-equilibrium kinetic calculation of N and O ionized mixtures, while our study is focused on monoatomic plasmas close to local thermodynamic equilibrium (LTE). However, their conclusions and the results in our work are qualitatively the same. Their conclusion is that the bulk viscosity coefficients are strongly affected by the electronic excitation which is quantitatively similar to ours for the influence of the ionization-recombination processes to the bulk viscosity and degree of ionization $\alpha$. Istomin and Kustova [12, Figure 3] in their study demonstrate that the bulk viscosity has a broad maximum as a function of the temperature. Moreover bulk viscosity significantly exceeds the shear one analogously to the atomic plasma which is the subject in the present study. For the hydrogen-helium cocktail in the present article we have derived for the dispersionless bulk viscosity $\zeta_{0}\propto(1-\alpha)\alpha$, and at fixed mass density $\rho$, the ionization degree $\alpha(T)$ monotonously increases between 0 and 1, and again as a function of the temperature we have a broad maximum.
Istomin & Kustova [12] emphasize how important is for calculation of bulk viscosity is to have reliable data for the temperature dependence of the rates of the chemical reactions. The main our motivation is that for a hydrogen like atom since more than 70 years we have a reliable formula for near threshold ionization cross=section which gives the opportunity to use analytical formula for the rate of ionization $\beta$.
In Ref [12] dense molecular plasma is studied, which arises in the Earth’s atmosphere at re-entrance of spacecrafts, for example. While the results of our work find application in the partially ionized solar chromosphere and its sound wave heating. Although these studies focus on different physical scenarios, high vs low dense plasma, the common conclusion is however the same, up to now the bulk viscosity has not been in the focus of plasma physics, despite many interesting applications. This is also noted in Ref. [12, Sec. Introduction]: However, in the majority of studies (except Refs. 10 and 12–14) atomic species are considered as structure-less particles. Under this assumption, no internal heat conductivity and bulk viscosity coefficients appear for atomic species, both neutral and ionized, which can cause the loss of accuracy in the transport terms evaluation.
II Saha equation and plasma thermodynamics. Cold plasmas approximation
II.1 Ideal gas approximation. System of notations
We analyze low density plasma for which its pressure is given by the ideal gas approximation
$p=n_{\mathrm{tot}}T,\quad n_{\mathrm{tot}}=$ $\mathcal{N}_{\mathrm{tot}}n_{\rho},$ (3)
$\mathcal{N}_{\mathrm{tot}}=\mathcal{N}_{a}+\mathcal{N}_{e},\quad n_{e}=\mathcal{N}_{e}\,n_{\rho},$ (4)
where $\mathcal{N}_{e}$ is the number of electrons per hydrogen atom and the electrons per unit volume is $n_{e}$ and $n_{\mathrm{tot}}$ is the total number of particles. Introducing atomic mass per hydrogen atom
$M^{*}=\sum_{a}\overline{a}_{a}M_{a},$ (5)
where $M_{a}$ is the atomic mass per atom a the mass density of the gas can be written as
$\rho=M^{*}n_{\rho}.$ (6)
It is convenient also to introduce averaged mass of the plasma particles including the electrons
$\langle M\rangle\equiv M^{*}/\mathcal{N}_{\mathrm{tot}}=\rho/n_{\mathrm{tot}},$ (7)
and the thermal velocity per averaged plasma particles
$c_{\mathrm{N}}^{2}=\frac{T}{\langle M\rangle}=\overline{p}/\overline{\rho}=\left(\frac{\partial p}{\partial\rho}\right)_{\!T}.$ (8)
As a rule, with an overline we denote thermally averaged quantities, meaning the system is in local thermodynamic equilibrium (LTE). The notations follow the same from Ref. [13].
II.2 The Saha equation
Now we can analyze the Saha equation for equilibrium ionization [14]. Each atom with atomic number $Z_{a}$ can be in a different degree of ionization which probability is denoted with $r_{i,a}$, where $i=0,\dots,Z_{a}$ shows the successive ionization state of the atom, i.e. $i=0$ is for the neutral state, $i=1$ is for the first ionized state and so on. In such a way the volume number of the corresponding ions
$n_{i,a}=n_{\rho}\overline{a}_{a}r_{i,a},\qquad\sum_{i=1}^{Z_{a}}\,r_{i,a}=1,$ (9)
and the charge neutrality condition gives the electron density
$n_{e}=n_{\rho}\,\mathcal{N}_{e},\qquad\mathcal{N}_{e}=\sum_{a}\overline{a}_{a}\sum_{i=1}^{Z_{a}}i\,r_{i,a}.$ (10)
Often the corresponding concentration is used [15, Sec. 104] and it is straightforward to change the notations
$c_{i,a}\equiv\frac{n_{i,a}}{n_{\mathrm{tot}}},\quad c_{e}\equiv\frac{n_{e}}{n_{\mathrm{tot}}},\quad n_{\mathrm{tot}}=n_{e}+n_{\rho}\sum_{a}\overline{a}_{a}.$ (11)
Within the framework of the up to now introduced notations, the Saha equation for the equilibrium concentrations reads in its standard form [15, Eq. (104.2)]
$\frac{\overline{c}_{i-1,a}}{\overline{c}_{i,a}\,\overline{c}_{e}}=\frac{\overline{p}}{T\tilde{n}_{i,a}},\qquad\overline{p}=\overline{n}_{\mathrm{tot}}\overline{T}=\overline{n}_{\mathrm{\rho}}\overline{\mathcal{N}}_{\mathrm{tot}}\overline{T},$ (12)
where following [13], we have introduced a chain of almost standard notations
$\tilde{n}_{i,a}\equiv\frac{g_{i,a}g_{e}}{g_{i-1,a}}n_{s,i,a},$ (13)
$n_{s,i,a}\equiv n_{q}\mathrm{e}^{-\iota_{i,a}},\quad\iota_{i,a}\equiv\frac{I_{i,a}}{T},\quad n_{q}\equiv\left(\frac{mT}{2\pi\hbar^{2}}\right)^{\!\!3/2}\!\!,$
$I_{i,a}$ is the $i$-th ionization potential of atom $a$, and $g_{i,a}$ are the statistical weights of the corresponding ions. Here we suppose that the plasma temperature is much higher than the fine energy splitting. For hydrogen, for example, $I_{1,\mathrm{H}}=13.6\,\mathrm{eV}$, $g_{1,\mathrm{H}}=2$, $g_{0,\mathrm{H}}=1$, and for the electrons we have two spin states and $g_{e}=2$.
It is numerically convenient to rewrite the Saha equation as
$\frac{\overline{r}_{i,a}}{\overline{r}_{i-1,a}}=\frac{\overline{n}_{i,a}}{\overline{n}_{i-1,a}}=\frac{\tilde{n}_{i,a}}{\overline{n}_{e}},$ (14)
and to express $\overline{r}_{i,a}$ by $\overline{r}_{i-1,a}$ starting with $r_{0,a}$ very small and $\overline{n}_{e}=\overline{n}_{\rho}/2$ [13]. Then we can make the normalization Eq. (9) and after calculating all $\overline{r}_{i,a}$, we can again calculate the electron concentration according to Eq. (10). This procedure is repeated until convergence with a required accuracy is reached. And last, we calculate the equilibrium pressure $\overline{p}$ using Eq. (12). Knowing the equilibrium degree of ionization, in the next subsection we calculate all thermodynamic variables.
II.3 Thermodynamic variables
At known ionization probabilities, we can easily calculate the ionization energy per atom for each atom
$E_{a}^{\mathrm{(ion)}}\equiv\sum_{i=1}^{Z_{a}}J_{i,a}r_{i,a},\qquad J_{j,a}\equiv\sum_{i=1}^{j}I_{i,a}.$ (15)
or in equilibrium
$\overline{E}_{a}^{\mathrm{(ion)}}\equiv\sum_{i=1}^{Z_{a}}J_{i,a}\overline{r}_{i,a}.$ (16)
Then in order to calculate the enthalpy per unit mass $w$, it is convenient to calculate the enthalpy per hydrogen atom
$M^{*}w=\mathcal{N}_{e}c_{p}T+\sum_{a}\overline{a}_{a}\left[c_{p}T+\sum_{i}r_{i,a}J_{i,a},\right],$ (17)
where $c_{p}=5/2$ is the heat capacity per atom at constant pressure. For the internal energy per hydrogen atom we have to substitute $c_{p}$ with the heat capacity per constant volume per one atom $c_{v}=3/2$
$M^{*}\varepsilon=$ $\,(\mathcal{N}_{e}+\mathcal{N}_{\mathrm{atom}})c_{v}T+\sum_{a}\overline{a}_{a}E_{a}^{\mathrm{(ion)}}$ (18)
$=$ $\sum_{a}\overline{a}_{a}\!\left[\left(1\!+\!\sum_{i=1}^{Z_{a}}ir_{i,a}\right)\!c_{v}T+\sum_{i=1}^{Z_{a}}r_{i,a}\sum_{j=1}^{i}I_{j,a}\right].$
For brevity, in equilibrium overline $\overline{r}_{i,a}$ in all those formulae have been omitted. In low frequency dynamics we can consider the time dependence $r_{i,a}(t)$ and temperature $T(t)$, and as a consequence: time dependent internal energy $\varepsilon(t)$ and enthalpy $w(t)$ per unit mass.
Knowing these thermodynamic potentials, we can easily calculate heat capacity per unit mass from the derivatives [15, Sec. 16]
$\mathcal{C}_{p}$ $\equiv\left(\dfrac{\partial w}{\partial T}\right)_{\!\!p},$ (19)
$\mathcal{C}_{v}$ $\equiv\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{\!\!\rho}=\left(\dfrac{\partial w}{\partial T}\right)_{\!\!\rho}-\dfrac{1}{\rho}\left(\dfrac{\partial p}{\partial T}\right)_{\!\!\rho}.$ (20)
Then the Jacobian [13, Eq. (9)]
$\mathcal{J}\equiv\frac{\partial(w,p)}{\partial(T,\rho)}=\left(\frac{\partial w}{\partial T}\right)_{\!\!\rho}\left(\frac{\partial p}{\partial\rho}\right)_{\!\!T}-\left(\frac{\partial w}{\partial\rho}\right)_{\!\!T}\left(\frac{\partial p}{\partial T}\right)_{\!\!\rho},$ (21)
can also be calculated by numerical differences if the enthalpy is calculated with almost machine accuracy.
The sound velocity at evanescent frequency can be calculated by the standard thermodynamic derivatives [13, Eq. (A9)]
$c_{0}^{2}(T,\rho)$ $=\left(\frac{\partial p}{\partial\rho}\right)_{\!\!s}=\frac{\mathcal{J}}{\mathcal{C}_{v}}$ (22)
$=\dfrac{\left(\dfrac{\partial w}{\partial T}\right)_{\!\!\rho}\left(\dfrac{\partial p}{\partial\rho}\right)_{\!\!T}-\left(\dfrac{\partial w}{\partial\rho}\right)_{\!\!T}\left(\dfrac{\partial p}{\partial T}\right)_{\!\!\rho}}{\left(\dfrac{\partial w}{\partial T}\right)_{\!\rho}-\dfrac{1}{\rho}\left(\dfrac{\partial p}{\partial T}\right)_{\!\!\rho}}.$
For our further analysis, this velocity is convenient to be represented by the generalized polytropic index
$\gamma_{0}$ $\equiv\gamma_{0}(T,\rho)\equiv\dfrac{\left(\dfrac{\partial p}{\partial\rho}\right)_{\!\!s}}{\left(\dfrac{\partial p}{\partial\rho}\right)_{\!T}}=\frac{c_{0}^{2}}{c_{\mathrm{N}}^{2}}=\frac{\rho}{p}\left(\frac{\partial p}{\partial\rho}\right)_{\!\!s}=\frac{\rho}{p}\cdot\dfrac{\mathcal{J}}{\mathcal{C}_{v}}$
$=\frac{\rho}{p}\cdot\dfrac{\left(\dfrac{\partial w}{\partial T}\right)_{\!\!\rho}\left(\dfrac{\partial p}{\partial\rho}\right)_{\!\!T}-\left(\dfrac{\partial w}{\partial\rho}\right)_{\!\!T}\left(\dfrac{\partial p}{\partial T}\right)_{\!\!\rho}}{\left(\dfrac{\partial w}{\partial T}\right)_{\!\rho}-\dfrac{1}{\rho}\left(\dfrac{\partial p}{\partial T}\right)_{\!\!\rho}}.$ (23)
In the formulae above, the overline denoting thermal equilibrium values have been omitted again. The velocity $c_{0}$ is the sound velocity at low frequency for which during the adiabatic compression of the plasma, the ionization degrees can follow the equilibrium Saha equation.
In the opposite limit when the ionization-recombination processes are negligible and the adiabatic compression is at constant chemical compound (isochema), for the sound velocity we have the standard monoatomic value
$c_{\infty}^{2}=\gamma_{\infty}\,c_{\mathrm{N}}^{2},\qquad\gamma_{\infty}=\gamma_{a}=5/3.$ (24)
Excluding these two limiting cases, the sound propagation with damping is determined by the kinetics of the ionization-recombination processes. For real frequencies, the wave amplitudes of the pressure and density are absorbed
$\delta p=\Re(\hat{p}(x,t)),\qquad\delta\hat{p}(x,t)\propto\exp(\mathrm{i}(\hat{k}x-\omega t)),$ (25)
$\delta\rho=\Re(\hat{\rho}(x,t)),\qquad\delta\hat{\rho}(x,t)\propto\exp(\mathrm{i}(\hat{k}x-\omega t)),$ (26)
$\mathrm{e}^{\mathrm{i}\hat{k}x}=\mathrm{e}^{-k^{\prime\prime}x}\mathrm{e}^{\mathrm{i}k^{\prime}x},\qquad\hat{k}=k^{\prime}+\mathrm{i}k^{\prime\prime},$ (27)
by the imaginary part of the frequency dependent wave number $k^{\prime\prime}(\omega)$.
The central purpose of the present study is to calculate the generalized polytropic index $\hat{\gamma}(\omega)$ describing the entropy production by the ionization-recombination processes parameterized by the complex phase velocity [16, Eq. (81.8)]
$\hat{c}_{\mathrm{phase}}^{2}\equiv\frac{\delta\hat{p}}{\delta\hat{\rho}}=\frac{\omega^{2}}{\hat{k}^{2}}=\hat{\gamma}(\omega)\,c_{\mathrm{N}}^{2}.$ (28)
In order to analyze these processes for cold plasmas, in the next section we consider in short the physics of near threshold ionization cross section for the simplest possible reaction of the ionization of a Hydrogen atom
$\rm e+H\rightarrow p+e+e.$ (29)
III Wannier ridge resonances and Wannier near threshold ionization cross-section
It is remarkable that the exact solution for the scattering in the Coulomb potential coincides with both Born and classical cross-section approximations. The physics is however, absolutely different. At high energies we have slightly perturbed plane waves while the classical approximation is applicable when the velocity $v$ of the scattering electron is much smaller than the Bohr velocity
$v_{\mathrm{{}_{B}}}\equiv\alpha_{\mathrm{{}_{S}}}\,c,\qquad\alpha_{\mathrm{{}_{S}}}\equiv\frac{e^{2}}{\hbar c},\qquad e^{2}\equiv\frac{q_{e}^{2}}{4\pi\varepsilon_{0}}.$ (30)
Let us analyze the quasi-classical motion of one $s$-electron around a nucleus. With zero $l=0$ angular momentum, the electron falls to the nucleus as a kamikaze comet to the Sun; but after banzai, it’s resurrected to a periodic motion.
Extremely interesting is the quasi-classical motion of two $s$-electrons. They move as mirror images with opposite radius vectors $\mathbf{r}_{1}=-\mathbf{r_{2}}$ and opposite momenta $\mathbf{p}_{1}=-\mathbf{p_{2}}$ like two electrons in a Cooper pair. This strongly correlated motion is an example of the inapplicability of the self-consistent Hartree-Fock approximation for this problem. However, strong correlations lead to another simplification: the effective Rydberg takes into account the mirror electron
$E_{n}=-\frac{\mathrm{R_{eff}}}{n^{2}},\qquad\mathrm{R_{eff}}=\frac{1}{2}\left(Z-\frac{1}{4}\right)mc^{2}\alpha_{\mathrm{{}_{S}}}^{2},$ (31)
where $Z$ is the charge of the rest ion and $m$ is the electron mass. The velocity of light $c$ disappears in all those formulae into the Sommerfeld fine splitting constant $\alpha_{\mathrm{{}_{S}}}$.
If we draw the effective potential of this 3-body system in some choice of variables, we observe a ridge. That is why those metastable resonances are called Wannier ridge resonances. and are observed by electron impact in noble gases [17, 18].
But how does an electron with small energy slightly above the threshold $I$ create the ionization reaction Eq. (29)? The two electrons escape quasi-classically with small energies and again the mirror images are important. This important problem was solved by Wannier in 1953 [17], see also [19, Sec. 147 Near-threshold cross section behavior], who obtained for the cross-section for the reaction Eq. (29)
$\sigma(\varepsilon)=C_{\mathrm{W}}a_{\mathrm{B}}^{2}\left(\frac{\varepsilon}{I}-1\right)^{\!\!\mathrm{w}}.$ (32)
The numerical values we use $\mathrm{w}\approx 1.18\sim 1$ and $C_{\mathrm{W}}\approx 2.7$ are obtained from the experimental study [20, Fig. 6]. We consider as a relevant discussion for plasma physics to explain that the Wannier formula is applicable within an acceptable accuracy for all $s$-electrons ionized by electron impact. The energy of the classical motion is just the quantum ionization potential.
Performing [21, Eq. (5)] the integration over equilibrium Maxwell distribution of electrons
$\beta=\langle v\sigma(\varepsilon)\rangle,\qquad v=\sqrt{2\varepsilon/m}.$ (33)
we obtain for the ionization rate of this impact ionization
$\beta(T)$ $\approx\frac{2}{\sqrt{\pi}}C_{\mathrm{W}}\Gamma(\mathrm{w}+1)\frac{\mathrm{e}^{-\iota}}{\iota^{\mathrm{w}-1/2}}\beta_{\mathrm{B}},$ (34)
$\beta_{\mathrm{B}}$ $\equiv v_{\mathrm{{}_{B}}}a_{\mathrm{{}_{B}}}^{2}=6.126\times 10^{-15}\,\mathrm{m^{3}/s}.$
The use of near-threshold cross-section is a good approximation only for cold plasmas with $T\ll I$ or $\iota\gg 1$. In many physical problems pure hydrogen plasma allows an analytical solution and for the ionization rate of cold hydrogen plasma we have an acceptable ab initio approximation [21].
IV Could plasmas kinetic equation and solution of linearized equation for harmonic pressure oscillations
IV.1 General kinetic equation
The cold plasma approximation $T\ll I_{a}$ or $\iota_{1,a}\gg 1$ can be easily checked by the solution of the Saha equation taking into account only the first ionized state of all elements in the cocktail $I_{a}\equiv I_{i=1,a}$ hence $\iota_{a}\equiv\iota_{i=1,a}$, and $\beta\rightarrow\beta_{a}$ in Eq. (34). We also have
$\sum_{i=2}^{Z_{a}}\overline{r}_{i,a}\ll 1$ (35)
and therefore it is convenient to re-denote
$\alpha_{a}\equiv r_{1,a}$ (36)
and to use the approximation
$r_{0,a}\approx 1-\alpha_{a},\quad r_{2,a}\approx 0,\quad r_{3,a}\approx 0,\quad\dots\,.$ (37)
For these relatively low temperatures we take into account only single electron ionization for the neutral atom
$\mathrm{e}+a_{0}\leftrightarrow a_{1}+\mathrm{e+e}.$ (38)
Our starting point is the general kinetic equation for the volume density of the $i$-th ionized state of the ion of atom $a$ $n_{i,a}(t)$
$\mathrm{d}_{t}n_{i,a}=n_{i,a}\frac{\mathrm{d}_{t}n_{\rho}}{n_{\rho}}$ $+\beta_{i,a}(T)\left[n_{e}n_{i-1,a}-\frac{n_{i,a}n_{e}^{2}}{\tilde{n}_{i,a}(T)}\right]$ (39)
$-\beta_{i+1,a}(T)\left[n_{e}n_{i,a}-\frac{n_{i+1,a}n_{e}^{2}}{\tilde{n}_{i+1,a}(T)}\right].$
where $\beta_{i,a}$ is the rate of single electron reaction
$\mathrm{a}^{i+}+\mathrm{e}\rightarrow\mathrm{a}^{(i+1)+}+2\mathrm{e}.$ (40)
We suppose that the multi-ionization processes have negligible contribution. Having taken the first ionized state, from Eq. (39) we have for $n_{1,a}(t)$
$\mathrm{d}_{t}n_{1,a}$ $=n_{1,a}\frac{\mathrm{d}_{t}n_{\rho}}{n_{\rho}}+\beta_{a}(T)\left[n_{e}n_{0,a}-\frac{n_{1,a}n_{e}^{2}}{\tilde{n}_{1,a}(T)}\right]$ (41)
with $\beta_{a}(T)$ given by Eq. (34) and Saha density parameter $\tilde{n}_{1,a}(T)$ given by Eq. (13). And for $n_{0,a}(t)$ from Eq. (36) and Eq. (37)
$\mathrm{d}_{t}n_{0,a}=-\mathrm{d}_{t}n_{1,a}.$ (42)
The temperature $T(t)$ and electron density
$n_{e}(t)\approx\sum_{a}n_{1,a}(t)=n_{\rho}(t)\sum_{a}\overline{a}_{a}\,r_{1,a}(t)$ (43)
and are also time $t$ dependent. The difference between Eq. (10) and the expression above Eq. (43) is that now we take into account only single ionization supposing that for cold plasmas higher ionizations are negligible.
IV.2 Linearization for small density oscillations
Our task is to study how small oscillations of the density
$\rho(t)=[1+\epsilon_{\rho}(t)]\,\overline{\rho},\quad\epsilon_{\rho}(t)=\Re(\hat{\epsilon}_{\rho}(t)),$ (44)
$\hat{\epsilon}_{\rho}(t)\propto\mathrm{e}^{-\mathrm{i}\omega t},\qquad|\hat{\epsilon}_{\rho}|\ll 1.$ (45)
create small relative oscillations of all variables around their respective equilibrium values
$T=\Re(\hat{T}(t)),\quad\;\;\hat{T}(t)=(1+\hat{\epsilon}_{{}_{T}})\,\overline{T},$ (46)
$p=\Re(\hat{p}(t)),\qquad\hat{p}(t)=(1+\hat{\epsilon}_{p})\,\overline{p},$ (47)
$p=\Re(\hat{p}(t)),\qquad\hat{p}(t)=(1+\hat{\epsilon}_{p})\,\overline{p},$ (48)
$\alpha_{a}=\Re(\hat{\alpha}_{a}(t)),\quad\hat{\alpha}_{a}(t)=(1+\hat{\epsilon}_{a})\,\overline{\alpha}_{a},$ (49)
$r_{i,a}=\Re(\hat{r}_{i,a}(t)),\;\;\hat{r}_{i,a}(t)=(1+\hat{\epsilon}_{i,a})\overline{r}_{i,a}.$ (50)
where all relative deviations $|\hat{\epsilon}|\ll 1$ are small and $\propto\mathrm{e}^{-\mathrm{i}\omega t}$. An alternative system of notations is to introduce the deviations
$\hat{T}=\overline{T}+\delta\hat{T},\qquad\hat{\epsilon}_{{}_{T}}=\delta\hat{T}/\overline{T},$ (51)
$\hat{p}=\overline{p}+\delta\hat{p},\qquad\;\;\hat{\epsilon}_{p}=\delta\hat{p}/\overline{p},$ (52)
$\hat{\rho}=\overline{\rho}+\delta\hat{\rho},\qquad\;\;\hat{\epsilon}_{\rho}=\delta\hat{\rho}/\overline{\rho}.$ (53)
It is convenient to work with linearized equations as complex numbers can be used and their real parts $\Re$ to be taken after the calculation.
In order to derive the linearized kinetic equation, we have to substitute these small deviations into the general kinetic equation Eq. (39) and finally in its cold plasma approximation Eq. (41). Omitting the complete derivation, we now mention only some technical details.
The obtained relations between the variations of the probabilities in the cold plasma approximation are
$r_{1,a}=\alpha_{a},\qquad r_{0,a}=1-\alpha_{a},\qquad r_{0,a}+r_{1,a}=1,$ (54)
$r_{1,a}=\overline{r}_{1,a}(1+\epsilon_{a}),\qquad r_{0,a}=\overline{r}_{0}(1+\epsilon_{0,a}),$ (55)
$\epsilon_{0,a}=-\frac{\overline{r}_{1,a}}{\overline{r}_{0,a}}\epsilon_{a}=-\frac{\overline{\alpha}_{a}}{1-\overline{\alpha}_{a}}\epsilon_{a},\quad\epsilon_{a}\equiv\epsilon_{1,a}=\Re{(\hat{\epsilon}_{a})}.$ (56)
Now we have to describe a chain of auxiliary notations: 1) Let us introduce
$\overline{X}_{a}=\beta_{a}(\overline{T})\,\overline{n}_{0.a}\,\overline{n}_{e}$ (57)
and take into account that
$\overline{n}_{0,a}=\overline{n}_{\rho}\,\overline{a}_{a}\,(1-\overline{\alpha}_{a}),\quad\overline{n}_{\rho}=\overline{\rho}/M^{*},\quad\overline{n}_{e}=\overline{n}_{\rho}\,\overline{\mathcal{N}}_{e}.$ (58)
2) As we suppose that density fluctuations are sinusoidal Eq. (44), for the time derivatives, following Heaviside, we introduce the imaginary variable
$\mathrm{D}=-\mathrm{i}\omega,\quad\mathrm{d}_{t}\hat{n}_{1,a}=\mathrm{D}\,\hat{n}_{1,a},\quad\mathrm{d}_{t}\hat{n}_{\rho}=\mathrm{D}\,\hat{n}_{\rho},\quad\dots\,.$ (59)
3) For the oscillations of the Saha densities $\tilde{n}_{1,a}(T(t))$ we have to take in the beginning that
$\delta\hat{T}^{\prime}\frac{\mathrm{d}\tilde{n}_{1,a}}{\mathrm{d}T}=(c_{v}+\iota_{1,a})\,\tilde{n}_{1,a}\hat{\epsilon}_{{}_{T}},\qquad\hat{\epsilon}_{{}_{T}}\equiv\frac{\delta\hat{T}}{\overline{T}}.$ (60)
4) For the calculations of the temperature oscillations $\delta\hat{T}$ we have to take into account the energy conservation. Let $\varepsilon$ be the internal energy per unit mass, we take the general expression which takes all possible ionization degrees Eq. (18).
For the general case, the set of kinetic equations for $n_{i,a}$ is completed by the equation from the temperature variation $\delta\hat{T}^{\prime}$ derived from the energy conservation of the liquid particle with unit mass
$\mathrm{d}\varepsilon=-p\,\mathrm{d}\mathcal{V},\qquad\mathcal{V}=1/\rho.$ (61)
Dividing of $\mathrm{d}\varepsilon$ from Eq. (18) by $\mathrm{d}t$ gives the differential equation
$\dot{\varepsilon}=-p\dot{\mathcal{V}}=\frac{p}{\rho^{2}}\,\dot{\rho}=\frac{p}{\rho}\,\mathrm{d}_{t}\epsilon_{\rho},\qquad\mathrm{d}_{t}=\frac{\mathrm{d}}{\mathrm{d}t}$ (62)
which writes as
$\frac{1}{M^{*}}\left[\mathcal{N}_{\mathrm{tot}}c_{v}\dot{T}+c_{v}T\dot{\mathcal{N}}_{e}+T\sum_{a}\overline{a}_{a}\sum_{i=1}^{Z_{a}}\overline{r}_{i,a}\dot{\epsilon}_{i,a}l_{i,a}\right],$
$=\frac{T}{M^{*}}\mathcal{N}_{\mathrm{tot}}\,\mathrm{d}_{t}\epsilon_{\rho},\qquad l_{i,a}\equiv J_{i,a}/T=\sum_{j=1}^{i}\iota_{i,a}.$ (63)
This linearized relation between first time derivatives gives for the small relative variations of the temperature; we have simply to omit $\mathrm{d}_{t}$ or over-dots with
$\epsilon_{{}_{T}}=\frac{\epsilon_{\rho}}{c_{v}}-\frac{\overline{\mathcal{N}}_{e}}{\overline{\mathcal{N}}_{\mathrm{tot}}}\epsilon_{e}-\frac{1}{c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}}\sum_{a}\overline{a}_{a}\sum_{i=1}^{Z_{a}}\overline{r}_{i,a}l_{i,a}\epsilon_{i,a},$ (64)
where the overline denoting thermodynamic equilibrium averaging can be omitted in the further linearized equations.
In such a way, from the general kinetic equation Eq. (39) we derive the linearized equation for the relative deviations from the equilibrium
$\hat{\epsilon}_{i,a}=\frac{\overline{n}_{e}}{\mathsf{D}}\left\{-\left(\beta_{i,a}\,\frac{\overline{r}_{i-1,a}}{\overline{r}_{i,a}}\right)(\hat{\epsilon}_{i,a}+\hat{\epsilon}_{e}-\hat{\epsilon}_{i-1,a})\right$. (65)
$\qquad\qquad+\beta_{i+1,a}\,(\hat{\epsilon}_{i+1,a}+\hat{\epsilon}_{e}-\hat{\epsilon}_{i,a})$
$\left.+\left[(c_{v}+\iota_{i,a})\left(\beta_{i,a}\,\frac{\overline{r}_{i-1,a}}{\overline{r}_{i,a}}\right)-(c_{v}+\iota_{i+1,a})\,\beta_{i+1,a}\right]\hat{\epsilon}_{{}_{T}}\right\}.$
Additionally the charge neutrality of the plasma gives for the electron density
$\mathcal{N}_{e}$ $=\frac{n_{e}}{n_{\rho}}=\sum_{a}\overline{a}\sum_{i=1}^{Z_{a}}i\overline{r}_{i,a}(1+\epsilon_{i,a}),$
$\dot{\mathcal{N}}_{e}$ $=\frac{n_{e}}{n_{\rho}}=\sum_{a}\overline{a}\sum_{i=1}^{Z_{a}}i\overline{r}_{i,a}\dot{\epsilon}_{i,a},$
$\mathcal{N}_{e}^{\prime}$ $=\int\dot{\mathcal{N}}_{e}\,\mathrm{d}t=\mathcal{N}_{e}-\overline{\mathcal{N}}_{e},$
$=\sum_{a}\overline{a}\sum_{i=1}^{Z_{a}}i\overline{r}_{i,a}\epsilon_{i,a}=\overline{\mathcal{N}}_{e}\,\epsilon_{e},$
$\hat{\epsilon}_{e}$ $=\frac{\hat{\mathcal{N}}_{e}^{\prime}}{\overline{\mathcal{N}}_{e}}=\frac{\sum_{a}\overline{a}\sum_{i=1}^{Z_{a}}i\overline{r}_{i,a}\hat{\epsilon}_{i,a}}{\sum_{a}\overline{a}\sum_{i=1}^{Z_{a}}i\overline{r}_{i,a}}$ (66)
One can introduce also relaxation rates having dimension of frequency
$f_{i,a}\equiv\frac{1}{\tau_{i,a}}\equiv\beta_{i,a}\overline{n}_{e},\qquad i\in(1,\,Z_{a}).$ (67)
For completeness we have to introduce ‘boundary conditions’ $f_{-1,a}=f_{Z_{a}+1,a}=0$. In such a way let us consider the physical meaning
$\tau_{i,a}^{(1/2)}\equiv\ln(2)\,\tau_{i,a}$ (68)
is the semi-decay period one ion a(i-1)+ to be ionized by electron impact end to become ai+.
One can trace that $\hat{\epsilon}_{{}_{T}}$ in Eq. (65) arises from the time dependent temperature in the Saha densities $\tilde{n}_{i,a}$ from Eq. (39). The substitution of $\hat{\epsilon}_{{}_{T}}$ from the equation of the temperature oscillations Eq. (64) into the linearized kinetic equation Eq. (65) for the relative oscillations of the concentrations gives and simultaneously $\hat{\epsilon}_{e}$ from Eq. (66) gives complete set of equations only for $\hat{\epsilon}_{i,a}$. We analyze this equation only in the cold plasma approximation.
IV.3 Numerical method for singly ionized cold plasmas
The general procedure for derivation of a set of linear equations for $\hat{\epsilon}_{i,a}$ we give in explicit form only in the cold plasma approximation which we study numerically. For the cold plasma approximation it is convenient to introduce the simplified notations
$\hat{\epsilon}_{a}\equiv\hat{\epsilon}_{1,a},\quad\hat{\epsilon}_{0,a}=-\frac{\overline{\alpha}_{a}}{1-\overline{\alpha}_{a}}\hat{\epsilon}_{a},$ (69)
$\hat{\epsilon}_{a,2}=\hat{\epsilon}_{a,3}=\dots=\hat{\epsilon}_{a,i\,>\,1}=0,$ (70)
$f_{a}\equiv\frac{1}{\tau_{a}}\equiv\beta_{1,a}\overline{n}_{e}.$ (71)
If we treat only ground and 1-st ionized state ($i\in[0,1]$), Eq. (10) gives
$\mathcal{N}_{e}$ $=\sum_{a}\overline{a}_{a}\sum_{i=1}^{Z_{a}}i\,r_{i,a}\approx\sum_{a}\overline{a}_{a}\,(1\,\,r_{1,a})=\sum_{a}\overline{a}_{a}\alpha_{a}$
$=\sum_{a}\overline{a}_{a}\overline{\alpha}_{a}(1+\epsilon_{1,a})=\overline{\mathcal{N}}_{e}(\epsilon_{e}+1),$
$\mathcal{\overline{N}}_{e}$ $=\sum_{a}\overline{a}_{a}\overline{\alpha}_{a}.$ (72)
This is valid in the case for low temperature plasma where $\iota_{1,a,1}\equiv I_{1,a}/T\gg 1$. The relation for the charge neutrality Eq. (66) now gives
$\hat{\epsilon}_{e}\approx\langle\hat{\epsilon}_{a}\rangle_{a}=\dfrac{\sum_{a}\overline{a}_{a}\overline{\alpha}_{a}\hat{\epsilon}_{a}}{\sum_{a}\overline{a}_{a}\overline{\alpha}_{a}}.$ (73)
Analogously general expression for the temperature oscillation, Eq. (64) takes the form
$\hat{\epsilon}_{{}_{T}}=\frac{\epsilon_{\rho}}{c_{v}}-\frac{\overline{\mathcal{N}}_{e}}{\overline{\mathcal{N}}_{\mathrm{tot}}}\hat{\epsilon}_{e}-\frac{1}{c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}}\sum_{a}\overline{a}_{a}\overline{\alpha}_{a}\iota_{a}\hat{\epsilon}_{a}$ (74)
Substituting here $\epsilon_{e}$ from Eq. (73) we obtain
$\hat{\epsilon}_{{}_{T}}=\frac{\epsilon_{\rho}}{c_{v}}-\frac{1}{c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}}\sum_{a}\overline{a}_{a}\overline{\alpha}_{a}(c_{v}+\iota_{a})\hat{\epsilon}_{a}.$ (75)
We substitute this relative temperature oscillation $\epsilon_{{}_{T}}$ and electron density oscillation $\hat{\epsilon}_{e}$ from Eq. (73) into the general kinetic equation Eq. (65) which for $\hat{\epsilon}_{1,a}$ reads
$\hat{\epsilon}_{a}=\dfrac{1-\overline{\alpha}_{a}}{1-\mathrm{i}\omega\tau_{a}\overline{\alpha}_{a}}\left\{\dfrac{\iota_{a}}{c_{v}}\hat{\epsilon}_{\rho}-\dfrac{1}{\overline{\mathcal{N}}_{\mathrm{tot}}}\left[\left(1+\dfrac{\iota_{a}}{c_{v}}\right)\sum_{b=1}^{a_{\mathrm{max}}}\iota_{b}\overline{a}_{b}\overline{\alpha}_{b}\hat{\epsilon}_{b}+\left(\iota_{a}+c_{v}+\dfrac{1}{\overline{c}_{e}}\right)\sum_{b=1}^{a_{\mathrm{max}}}\overline{a}_{b}\overline{\alpha}_{b}\hat{\epsilon}_{b}\right]\right\}.$ (76)
Here for each $a=$ H, He, C, Mg, Si, Fe we have to solve the corresponding linear equation for $a_{\mathrm{max}}=6$ variables $\hat{\epsilon}_{a}$. In order to alleviate numerical programming we introduce some convenient notations
$A_{a}=\overline{a}_{a}\overline{\alpha}_{a},\qquad B_{a}=\iota_{a}\overline{a}_{a}\overline{\alpha}_{a}=\iota_{a}A_{a},$ (77)
$\hat{\Gamma}_{a}=\dfrac{1-\overline{\alpha}_{a}}{1-\mathrm{i}\overline{\alpha}_{a}\tau_{a}},\qquad a=1,\,\ 2,\,\dots a_{\mathrm{max}},$ (78)
$C_{a}=\left(\iota_{a}+c_{v}+1/\overline{c}_{e}\right)/\overline{\mathcal{N}}_{\mathrm{tot}},$ (79)
$D_{a}=\left(1+\iota_{a}/c_{v}\right)/\overline{\mathcal{N}}_{\mathrm{tot}},$ (80)
the row the matrices
$\sf{L}_{A}=\begin{pmatrix}A_{1}&A_{2}&\dots&A_{a_{\mathrm{max}}}\\ \end{pmatrix}$ (81)
$\sf{L}_{B}=\begin{pmatrix}B_{1}&B_{2}&\dots&B_{a_{\mathrm{max}}}\end{pmatrix}$ (82)
and the square matrices $a_{\mathrm{max}}\times a_{\mathrm{max}}$,
$\hat{\sf{M}}_{A}=\begin{pmatrix}\hat{\Gamma}_{1}C_{1}\sf{L}_{A}\\ \hat{\Gamma}_{2}C_{2}\sf{L}_{A}\\ \dots\\ \hat{\Gamma}_{a_{\mathrm{max}}}C_{a_{\mathrm{max}}}\sf{L}_{A}\\ \end{pmatrix},\qquad\hat{\sf{M}}_{B}=\begin{pmatrix}\hat{\Gamma}_{1}D_{1}\sf{L}_{B}\\ \hat{\Gamma}_{2}D_{2}\sf{L}_{B}\\ \dots\\ D_{a_{\mathrm{max}}}\sf{L}_{B}\\ \end{pmatrix}$ (83)
and
[…]
and the column matrices
$\hat{\sf{R}}=\begin{pmatrix}\hat{\Gamma}_{1}\iota_{1}\\ \hat{\Gamma}_{2}\iota_{2}\\ \dots\\ \hat{\Gamma}_{a_{\mathrm{max}}}\iota_{a_{\mathrm{max}}}\\ \end{pmatrix}\frac{\epsilon_{\rho}}{c_{v}},\qquad\hat{\epsilon}=\begin{pmatrix}\hat{\epsilon}_{1}\\ \hat{\epsilon}_{2}\\ \dots\\ \hat{\epsilon}_{a_{\mathrm{max}}}\\ \end{pmatrix}$. (85)
Now the solution of Eq. (76) reads
$\hat{\sf{M}}\hat{\epsilon}=\hat{\sf{R}},\qquad\hat{\epsilon}=\hat{\sf{M}}^{-1}\hat{\sf{R}}$ (86)
and Eq. (73) and Eq. (75) can be rewritten as
$\hat{\epsilon}_{e}=\sf{L}_{A}\cdot\hat{\epsilon}/\overline{\mathcal{N}}_{e},$ (87)
$\hat{\epsilon}_{{}_{T}}=\hat{\epsilon}_{\rho}/c_{v}+\overline{c}_{e}\hat{\epsilon}_{e}+\sf{L}_{B}\cdot\hat{\epsilon}/\overline{\mathcal{N}}_{\mathrm{tot}},$ (88)
where the central dot $\cdot$” denotes matrix multiplication.
After the numerical solution of the kinetic equation we calculate all physical variables. The density deviation from equilibrium from Eq. (4)
$n_{\mathrm{tot}}=\overline{n}_{\mathrm{tot}}(1+\epsilon_{n})=n_{\rho}\mathcal{N}_{\mathrm{tot}}$ (89)
is easily derived in terms of the electron and mass density deviations
$\epsilon_{n}=\epsilon_{\rho}+\overline{c}_{e}\,\epsilon_{e},$ (90)
where the ratio $\overline{c}_{e}\equiv\overline{\mathcal{N}}_{e}/\overline{\mathcal{N}}_{\mathrm{tot}}\in(0,\,1)$ is the electron concentration in equilibrium. Analogously for the pressure from Eq. (3)
$p=\overline{p}\,(1+\epsilon_{p})=n_{\mathrm{tot}}\,T=\overline{n}_{\mathrm{tot}}\,(1+\epsilon_{n})\,T_{0}\,(1+\epsilon_{{}_{T}})$ (91)
and its deviation from equilibrium
$\hat{\epsilon}_{p}=\hat{\epsilon}_{n}+\hat{\epsilon}_{{}_{T}}=\epsilon_{\rho}+\hat{\epsilon}_{{}_{T}}+\overline{c}_{e}\hat{\epsilon}_{e},$ (92)
where
$\rho=M^{*}n_{\rho}=\overline{\rho}\,(1+\epsilon_{\rho}),\qquad M^{*}\equiv\sum_{a}\overline{a}_{a}M_{a}.$ (93)
After this kinetic analysis, we can address to the initial hydrodynamic problem of sound propagation in homogeneous cold plasmas.
IV.4 Generalized polytropic index for cold plasmas
Having calculated all necessary deviations from equilibrium, we can further proceed our analysis with the sound waves absorption. For that purpose following [16, Sec. 81, Second viscosity] at real frequency $\omega$ we can introduce complex wave vector $\hat{k}=k^{\prime}+\mathrm{i}k^{\prime\prime}$ and complex phase velocity
$\hat{c}^{2}_{\mathrm{phase}}$ $\equiv\frac{\omega^{2}}{\hat{k}^{2}(\omega)}\equiv\left(\frac{\partial p}{\partial\rho}\right)_{\!\!\omega}\equiv\frac{\delta\hat{p}}{\delta\hat{\rho}}=c_{\mathrm{{}_{N}}}^{2}\,\hat{\gamma}(\omega),$ (94)
$\hat{\gamma}(\omega)$ $\equiv\frac{\hat{\epsilon}_{p}}{\hat{\epsilon}_{\rho}}=1+\frac{\hat{\epsilon}_{{}_{T}}+\overline{c}_{e}\hat{\epsilon}_{e}}{\hat{\epsilon}_{\rho}}$ (95)
$=\gamma_{a}-\frac{1}{c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}\hat{\epsilon}_{\rho}}\sum_{a=1}^{a_{\mathrm{max}}}\overline{a}_{a}\overline{\alpha}_{a}\iota_{a}\hat{\epsilon}_{a}$ (96)
$=\gamma_{a}-\frac{\sf{L}_{B}\cdot\hat{\epsilon}}{c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}\hat{\epsilon}_{\rho}}=\gamma^{\prime}+\mathrm{i}\gamma^{\prime\prime},$ (97)
where $c_{\mathrm{{}_{N}}}^{2}=\overline{p}/\overline{\rho}.$
As test for the working programs we have to check that for very low and very high frequencies the polytropic index is real only. For high frequencies it reaches the atomic value $\gamma_{a}$ corresponding to negligible chemical reactions. For low frequencies we have a sophisticated thermodynamic expression Eq. (23) [13, Eq. (8)]
$\hat{\gamma}(\omega\rightarrow\infty)=\gamma_{a}\equiv 5/3,\qquad\gamma^{\prime\prime}(\infty)\rightarrow 0,\quad\gamma^{\prime}(\infty)\rightarrow\gamma_{a},$
$\hat{\gamma}(\omega\rightarrow 0)\approx\gamma^{\prime}(0)=\gamma_{0},\qquad\gamma^{\prime\prime}(0)\rightarrow 0.$ (98)
For the linear analysis in the programs we can choose $\epsilon_{\rho}=1$. For some purposes it is instructive also to introduce relative “polytropic” index
$\hat{\gamma}_{\mathrm{rel}}\equiv\dfrac{\hat{\gamma}(\omega)}{\gamma_{\infty}}=1-\frac{\sf{L}_{B}\cdot\hat{\epsilon}}{c_{p}\,\overline{\mathcal{N}}_{\mathrm{tot}}\,\hat{\epsilon}_{\rho}},$ (99)
and for the complex “phase” velocity we now have
$\hat{c}^{2}_{\mathrm{phase}}=\frac{\omega^{2}}{\hat{k}^{2}(\omega)}=\hat{\gamma}_{\mathrm{rel}}(\omega)\,c_{\infty}^{2}.$ (100)
Having an explicit expression for $\hat{\gamma}_{\mathrm{rel}}(\omega)$, we can calculate the complex wave-vector
$\hat{k}=\frac{\omega}{c_{\infty}\sqrt{\hat{\gamma}_{\mathrm{rel}}(\omega)}}=\frac{\omega}{c_{\mathrm{N}}\sqrt{\hat{\gamma}(\omega)}}=k^{\prime}+\mathrm{i}k^{\prime\prime}$ (101)
and its imaginary part gives the sound wave damping created by the bulk viscosity.
V Analysis of the results derived by the kinetic equation
V.1 Mandelstam-Leontovich and Cole-Cole approximation
We have solved the low temperature kinetic equation Eq. (86) for the solar plasma and calculated the dimensionless ratio
$\Gamma_{0,\infty}\equiv\frac{\gamma_{0}}{\gamma_{\infty}}=\dfrac{c_{0}^{2}}{c_{\infty}^{2}}=0.654<1.$ (102)
The calculated frequency dependence $\hat{\gamma}(\omega)=\gamma^{\prime}+\mathrm{i}\gamma^{\prime\prime}$ deserves to be compared with the Mandelstam-Leontovich (ML) single time-constant approximation cf. [16, Eq. 81.8]
$\hat{\gamma}_{\mathrm{rel}}^{\mathrm{(ML)}}(\omega)=\dfrac{\Gamma_{0,\infty}-\mathrm{i}\omega\tau}{1-\mathrm{i}\omega\tau}=\dfrac{\Gamma_{0,\infty}+\tilde{s}\tau}{1+\tilde{s}\tau},$ (103)
where $\tilde{s}\equiv\mathrm{i}\omega$. The comparison is represented in the Argand plot $\gamma^{\prime}$ versus $\gamma^{\prime\prime}$ drawn in Fig. 1. We were surprised that the curves practically coincide.
[…]
The parameter $\Gamma_{0,\infty}$ related with the parameter $\mathcal{M}_{\mathrm{RS}}$ used by Rudenko and Soluyan [23] can be also expressed as
$\Gamma_{0,\infty}\equiv\ \frac{1}{1+\mathcal{M}_{\mathrm{RS}}},\qquad\mathcal{M}_{\mathrm{RS}}\equiv\frac{c_{\infty}^{2}}{c_{0}^{2}}-1$ (104)
see also [21, Eq. 99].
The excellent accuracy of the ML fit with fitting parameter time-constant $\tau$ gives a simple interpretation of the both low- and high-frequency limits; we have a simple Padé approximation. We can consider the oscillation of the density as a perturbation and the pressure as a linear response. According to the causality principle, the generalized susceptibility $\hat{p}/\hat{\rho}\propto\hat{\gamma}(\omega)$ must be analytical in the upper semi-plane of the complex frequency $\Im(\omega)>0$ and just along the imaginary semi-axis $\omega=\mathrm{i}\tilde{s}$ with $\tilde{s}>0$ the compressibility must be real and positive. In the lower imaginary semi-plane $\hat{\gamma}_{\mathrm{rel}}^{\mathrm{{ML}}}(\omega)$ has a zero at $\omega=-\mathrm{i}/\tau$ and a pole at $\omega=-\mathrm{i}\Gamma_{0,\infty}/\tau$ in agreement with the general causality principle [15, Sec. 123 Generalized susceptibility].
One can introduce also Q-factor for the sound waves damped by bulk viscosity
$Q_{a}(\omega)\equiv\frac{k^{\prime}}{2k^{\prime\prime}}=\frac{\Re\left(1/\sqrt{\hat{\gamma}_{\mathrm{rel}}(\omega)}\right)}{\Im\left(2/\sqrt{\hat{\gamma}_{\mathrm{rel}}(\omega)}\right)}\gg 1.$ (105)
This factor is always high $Q_{a}=k^{\prime}/2k^{\prime\prime}\gg 1$. One can define the real sound velocity [16, Sec. 81, Second viscosity]
$c_{\mathrm{s}}(\omega)\equiv c_{\infty}\Re\left(\sqrt{\hat{\gamma}_{\mathrm{rel}}}\right)=c_{\mathrm{{}_{N}}}\Re\left(\sqrt{\hat{\gamma}}\right)\approx\Re(\hat{c}_{\mathrm{phase}}).$ (106)
For the large enough values of the Q-factor, one can represent relative polytropic index
$\hat{\gamma}_{\mathrm{rel}}(\omega)=\gamma_{\mathrm{rel}}^{\prime}+\mathrm{i}\gamma_{\mathrm{rel}}^{\prime\prime}=|\hat{\gamma}_{\mathrm{rel}}|\mathrm{e}^{\mathrm{i}\varphi_{\gamma}}\approx\gamma_{\mathrm{rel}}^{\prime}(1+\mathrm{i}\varphi_{\gamma}),$ (107)
$\hat{k}=k^{\prime}+\mathrm{i}k^{\prime\prime}=|k|\mathrm{e}^{\mathrm{i}\varphi_{k}}\approx(1+\mathrm{i}/2Q_{a})k^{\prime},$ (108)
$|\varphi_{\gamma}|\ll 1,\qquad\varphi_{k}=-\frac{1}{2}\varphi_{\gamma}=\frac{1}{2Q_{a}}.$ (109)
The general formula for the wave-vector frequency dependence of the complex wave-vector, cf. Ref. [16, Eq. (81.9-10)]
$\hat{k}\equiv$ $k^{\prime}+\mathrm{i}k^{\prime\prime}\approx\frac{\omega}{c_{\mathrm{s}}}+\mathrm{i}\frac{\omega^{2}\zeta^{\prime}(\omega)}{2\rho_{0}c_{\mathrm{s}}^{3}}$ (110)
$=$ $\left(1+\mathrm{i}\frac{\omega\zeta^{\prime}}{2\rho c_{\mathrm{s}}^{2}}\right)\frac{\omega}{c_{\mathrm{s}}}=\left(1-\mathrm{i}\frac{\varphi_{\gamma}}{2}\right)k^{\prime}$ (111)
gives the low frequency limit [16, Eq. (81.9)] for the ML approximation Eq. (103)
$\hat{k}(\omega\ll 1/\tau)\approx\frac{\omega}{c_{0}}\left[1+\mathrm{i}\omega\tau\left(\frac{1}{\Gamma_{0,\infty}}-1\right)\right]$ (112)
and correspondingly high-frequency approximation [16, Eq. (81.10)] ML approximation Eq. (103)
$\hat{k}(\omega\gg 1/\tau)=k^{\prime}+\mathrm{i}k^{\prime\prime}\approx\frac{\omega}{c_{\infty}}\left[1+\mathrm{i}\left(1-\Gamma_{0,\infty}\right)\right].$ (113)
The frequency dependent reciprocal Q-factor of the ML approximation Eq. (103)
$\frac{1}{Q_{a}^{\mathrm{(ML)}}(\omega)}\equiv-\frac{2\Im\left(\sqrt{\hat{\gamma}_{\mathrm{rel}}^{\mathrm{(ML)}}(\omega)}\right)}{\Re\left(\sqrt{\hat{\gamma}_{\mathrm{rel}}^{\mathrm{(ML)}}(\omega)}\right)}$ (114)
is depicted in Fig. 2
Figure 2: Reciprocal Q-factor of the ML approximation as a function of $x\equiv\omega\tau$ according to Eq. (114) and Eq. (103) $1/Q_{a,\mathrm{max}}\equiv Q_{a,\mathrm{min}}=0.21$ at $x_{\mathrm{max}}=0.8235$. Independently using Eq. (105), we obtain $\omega_{\mathrm{max}}=0.9127$ µrad/s and the corresponding time-constant is determined as $\tau_{\mathrm{max}}=x_{\mathrm{max}}/\omega_{\mathrm{max}}\approx 10.5$ days. In the general case $1/\mathcal{Q}^{\mathrm{(ML)}}$ has a maximum at $x_{\mathrm{max}}=\omega_{\mathrm{max}}\tau_{\mathrm{max}}=\sqrt{\Gamma_{0,\infty}}=c_{0}/c_{\infty}$, cf. [16, Sec. 81]. Here we have plotted only the $1/\mathcal{Q}^{\mathrm{(ML)}}_{a}(\omega)$ from Eq. (114) since the same calculation from Eq. (105) coincides with it as the Cole-Cole plot in Fig. 1
This function $\left[Q_{a}^{\mathrm{(ML)}}\right]^{-1}$ has maximum at argument $x\equiv\omega\tau=\sqrt{\Gamma_{0,\infty}},$ where the parameter $\Gamma_{0,\infty}=c_{0}^{2}/c_{\infty}^{2}=0.654$ is determined by the ML fit represented in Fig. 1.
[…]
$\tau_{\mathrm{max}}=x_{\mathrm{max}}/\omega_{\mathrm{max}}$ (115)
is calculated. The exact value of $\tau_{\mathrm{max}}$ is irrelevant for the Argand plot in Fig. 1 where the exact complex function $\hat{\gamma}$ and its ML approximation $\hat{\gamma}^{\mathrm{(ML)}}$ are represented.
The frequency independent height-frequency asymptotic gives the possibility to calculate in short wavelength approximation the absorption coefficient
$A=\exp\left(-\int_{0}^{h_{\mathrm{TR}}}2k^{\prime\prime}(x)\mathrm{d}x\right)$ (116)
and to evaluate the total heating power induced to the chromosphere by the high frequency part of the energy flux of the acoustic waves
$\int_{0}^{h_{\mathrm{TR}}}\mathcal{Q}_{\zeta}(x)\,\mathrm{d}x=A\,q_{\mathrm{ac}}.$ (117)
Taking in Eq. (18) only singe ionization we obtain
$\overline{\varepsilon}=\frac{T}{M^{*}}\left(\overline{\mathcal{N}}_{\mathrm{tot}}c_{v}+\sum_{a}\overline{a}_{a}\overline{\alpha}_{a}\iota_{a}\right),$ (118)
$\overline{w}=\frac{T}{M^{*}}\left[\overline{\mathcal{N}}_{\mathrm{tot}}c_{p}+\sum_{a}\overline{a}_{a}\overline{\alpha}_{a}\iota_{a}\right]$ (119)
where we repeat the definitions for the number or atoms and equilibrium electrons per hydrogen atom
$\overline{\mathcal{N}}_{\mathrm{tot}}=\overline{\mathcal{N}}_{a}+\overline{\mathcal{N}}_{e},\quad\overline{\mathcal{N}}_{a}=\sum_{a}\overline{a}_{a},\quad\overline{\mathcal{N}}_{e}=\sum_{a}\overline{a}_{a}\overline{\alpha}_{a}.$
V.2 General formula for the bulk viscosity expressed by complex polytropic index $\hat{\gamma}(\omega)$ and Drude approximation for $\hat{\zeta}(\omega)$
Let us write again the formula for the wave vector with small imaginary part created by bulk viscosity Eq. (110), confer Ref. [16, Eq. (79.6) and Eq. (81.9-10)]
$\hat{k}(\omega)$ $=k^{\prime}+\mathrm{i}k^{\prime\prime}=\left[1+\mathrm{i}\frac{\omega\,\zeta^{\prime}(\omega)}{2\,\overline{\rho}\,c_{\mathrm{s}}^{2}(\omega)}\right]\frac{\omega}{c_{\mathrm{s}}(\omega)}$ (120)
$=\frac{\omega}{c_{\mathrm{N}}\sqrt{\hat{\gamma}(\omega)}}\approx\left[1-\mathrm{i}\frac{\gamma^{\prime\prime}}{2\gamma^{\prime}}\right]\frac{\omega}{c_{\mathrm{s}}},$ (121)
$k^{\prime}(\omega)$ $=\Re(\hat{k}),\quad k^{\prime\prime}(\omega)=\Im(\hat{k}),\quad c_{\mathrm{s}}\equiv c_{\mathrm{{}_{N}}}\sqrt{\gamma^{\prime}}.$ (122)
For $|\gamma^{\prime\prime}|\ll\gamma^{\prime}$ the comparison gives the general formula for the bulk viscosity expressed by the imaginary part of polytropic index
$\zeta^{\prime}(\omega)=-\frac{\gamma^{\prime\prime}(\omega)}{\omega}\,\overline{p},\qquad\frac{\zeta^{\prime}(\omega)}{\overline{\rho}}=-\frac{\gamma^{\prime\prime}(\omega)}{\omega}\,c_{\mathrm{{}_{N}}}^{2}.$ (123)
which is one of the main results of the present study.
The sound velocity ratio $c_{\mathrm{s}}(\omega)/c_{\mathrm{{}_{N}}}=\sqrt{\gamma^{\prime}}$ as function of frequency $f=\omega/2\pi$ is depicted at Fig. 3.
Figure 3: Sound velocity ration $c_{\mathrm{s}}/c_{\mathrm{{}_{N}}}=\sqrt{\gamma^{\prime}}$ as function of frequency $f=\omega/2\pi$ according to the calculations by the general formula Eq. (97). For high frequencies $\sqrt{\gamma^{\prime}(\infty)}=\sqrt{5/3}=\sqrt{\gamma_{a}}$ (upper horizontal dashed line) this ratio reaches mono-atomic value. At low frequency $\sqrt{\gamma^{\prime}(0)}=\sqrt{\gamma(T,\rho)}\simeq 1$ sound velocity is determined by the explicit thermodynamic expression Eq. (23). The minimal value $\gamma_{0}\approx 1$ (the lower horizontal dashed line) can be reached at $T\ll I_{1}$ for $\overline{\alpha}_{1}\simeq 1/2$.
The ratio with dimension of time
$\frac{\zeta^{\prime}(\omega)}{\overline{p}}=-\frac{\gamma^{\prime\prime}(\omega)}{\omega}$ (124)
is shown in Fig. 4.
Figure 4: Frequency dependence of the real (absorptive) part of the bulk viscosity according to Eq. (124). On the ordinate is drawn $\zeta^{\prime}/\overline{p}$ with dimension of time, while on the abscissa is the angular frequency $\omega$.
The low frequency $\zeta_{0}$ limit has to be calculated for $\omega\tau\ll 1$; formally
$\zeta_{0}=-\overline{p}\lim_{\omega\rightarrow 0}\frac{\gamma^{\prime\prime}(\omega)}{\omega}=-\overline{p}\left.\frac{\mathrm{d}\gamma^{\prime\prime}}{\mathrm{d}\omega}\right|_{\omega=0}.$ (125)
For very low frequencies for $\gamma_{0}=\gamma(\omega=0)$ we have the explicit thermodynamical expression by the Jacobian Eq. (23). For a cocktail of alkali metals and noble gases we even have an explicit solution Eq. (178). In such a way, the frequency dependent kinetic coefficient
$\hat{\zeta}(\omega)=\mathrm{i}\,\frac{\hat{\gamma}(\omega)}{\omega}\,\overline{p}$ (126)
is obtained by the solution of the kinetic equation applied to the pressure when we consider density fluctuations as a driving perturbation.
The time-constant $\tau$ from the good working (see Fig. 1) Mandelstam-Leontovich approximation Eq. (103) can be evaluated using [16, Eq. (81.7)]
$\zeta_{0}=\tau\overline{\rho}(c_{\infty}^{2}-c_{0}^{2}),$ (127)
which in our notations can be rewritten as
$\zeta_{0}=\tau\,\overline{p}\,(\gamma_{\infty}-\gamma_{0}).$ (128)
At very small and very high frequencies the imaginary part of the polytropic index is zero
$\gamma_{0}^{\prime\prime}=\gamma_{\infty}^{\prime\prime}=0,\qquad\gamma_{0}^{\prime}=\gamma_{0}=\gamma(T,\rho),\quad\gamma_{\infty}^{\prime}=\gamma_{a}.$ (129)
The comparison of this interpolation expression with the low-frequency result gives the final result for the Mandelstam-Leontovich time constant expressed by the complex frequency depended polytropic index
$\tau_{0}\equiv-\frac{1}{\gamma_{\infty}^{\prime}-\gamma_{0}^{\prime}}\lim_{\omega\rightarrow 0}\frac{\gamma^{\prime\prime}(\omega)}{\omega},$ (130)
$\frac{\gamma_{\infty}^{\prime}-\gamma_{0}^{\prime}}{1/\tau}=-\left.\frac{\mathrm{d}\gamma^{\prime\prime}(\omega)}{\mathrm{d}\omega}\right|_{\omega=0},$ (131)
where $\hat{\gamma}(\omega)=\gamma^{\prime}+\mathrm{i}\gamma^{\prime}$ is given by the sum Eq. (97) over small variations of the ionization degrees $\hat{\epsilon}_{a}$ created by the density oscillations $\hat{\epsilon}_{\rho}\propto\mathrm{e}^{-\mathrm{i}\omega t}$. The time constant $\tau_{0}\approx 10.25$ days according to Eq. (130) which compared with the Mandelstam-Leontovich fitted $\tau_{\mathrm{max}}\approx 10.5$ days from Eq. (114) and in Fig. 2 has $\approx 2\%$ difference. This agreement is a consequence of the good working ML fit plotted in Fig. 1.
Using the so determined $\tau$, one can evaluate [16, Eq. (81.9-10)]
$k^{\prime\prime}(\omega\rightarrow 0)=\frac{(c_{\infty}^{2}-c_{0}^{2})\tau\omega^{2}}{2c_{0}^{3}}=\frac{\zeta_{0}}{2\rho c_{0}^{3}}\,\omega^{2},$ (132)
$k^{\prime\prime}(\omega\rightarrow\infty)=\frac{c_{\infty}^{2}-c_{0}^{2}}{2\tau c_{\infty}^{3}}=\frac{\zeta_{0}}{2\tau^{2}\rho c_{\infty}^{3}}=\mathrm{const},$ (133)
and also Drude frequency dependence of the complex bulk viscosity [16, Eq. (81.6)]
$\hat{\zeta}(\omega)=\zeta^{\prime}+\mathrm{i}\zeta^{\prime\prime}\approx\frac{\zeta_{0}}{1-\mathrm{i}\omega\tau},\quad\zeta^{\prime}(\omega)\approx\frac{\zeta_{0}}{1+\omega^{2}\tau^{2}}.$ (134)
In Fig. 5 are depicted frequency dependent bulk viscosity and its Mandelstam-Leontovich approximation.
$\mathcal{Z}\equiv\zeta^{\prime}(\omega)/\zeta_{0},\qquad\mathcal{W}=\omega\tau,\qquad 1/\mathcal{Z}\approx 1+\mathcal{W}^{2}$ (135)
obtained by solution of the kinetic equation Eq. (76).
Figure 5: Frequency dependent dimensionless bulk viscosity $\mathcal{Z}=\zeta^{\prime}/\zeta_{0}$ as function of dimensionless frequency $\mathcal{W}\equiv\omega\tau$ according according to Eq. (76). The curve is undistinguished by its Drude fit.
We consider that this plot will be useful for future experimental data processing.
Having reliable frequency dependence of the absorption part of the bulk viscosity Eq. (134) with the routine calculation of $\hat{\zeta}(\omega)$ Eq. (97), we have an alternative method for determination of the ML relaxation time
$\tau_{\infty}\equiv\lim_{\omega\rightarrow\infty}\left(\,\frac{1}{\omega}\sqrt{\frac{\zeta_{0}}{\zeta^{\prime}(\omega)}-1}\,\right)\approx\sqrt{\frac{\zeta_{0}}{\omega^{2}\zeta^{\prime}(\omega)}}.$ (136)
The numerical agreement of the so defined $\tau_{\infty}$ Eq. (136) with $\tau_{0}$ Eq. (130) and $\tau_{\mathrm{max}}$ from Eq. (115) is the crucial test for the applicability of the ML approximation. Agreement within several percent accuracy is a simple consequence that at low temperature plasma, the number of a single type ions dominates. In this case, imagine pure hydrogen, we have only one equation in Eq. (76) for the relative amplitude of oscillation hydrogen ionization degree $\hat{\epsilon}_{1}$. Substitution in all general formulae reveals that for pure hydrogen, plasma ML approximation is actually the exact solution.
That is why the so determined time-constant $\tau$ can be used even for the high frequency case $\omega\tau\gg 1$ as it is represented in Fig. 6.
Figure 6: The same results from the exact numerical solution of the kinetic equation Eq. (76) represented in Fig. 5 but in the plot: $1/\mathcal{Z}$ in the ordinate versus $\mathcal{W}^{2}$ in the abscissa. For us it was the most complicated method to draw with the computer pixel accuracy the straight line $1/\mathcal{Z}=1+\mathcal{W}^{2}$ of the ML approximation. The final result in the present study is that the ML approximation has an excellent accuracy for cold plasmas with $T\ll I_{a}$.
This redrawing in Fig. 6 of the Drude behavior as a straight line according to Eq. (135) was done in order to observe any deviations of the calculated $\zeta(\omega)$ from its ML approximation $\zeta^{\mathrm{ML}}(\omega)$. We did not expect that this approximation will be so good.
The dimensionless complex function $\hat{\gamma}(\omega)=\gamma^{\prime}+\mathrm{i}\gamma^{\prime\prime}$ is the most important ingredient of our theoretical analysis of the bulk viscosity of the cold plasma. In Fig. 7 its real and imaginary part are depicted.
Figure 7: Frequency dependence of the complex adiabatic index $\hat{\gamma}=\gamma^{\prime}+\mathrm{i}\gamma^{\prime\prime}$ as a function of the dimensionless frequency $\mathcal{W}=\omega\tau$. These functions together with the Argand plot Fig. 1 $\gamma^{\prime}$ versus $\gamma^{\prime\prime}$ gives the complete function $\hat{\gamma}(\omega\tau)$. The real part $\gamma^{\prime}$ monotonously increases from $\gamma_{0}$ to $\gamma_{a}=5/3$, while the absorptive part $\gamma^{\prime\prime}$ has a maximum at $\mathcal{W}_{\mathrm{max}}=\omega_{\mathrm{max}}\tau\approx 1$, confer Fig. 2: the accuracy of this approximation depends of the accuracy of ML fit according Fig. 1. Its tangent at zero frequency $\omega[-\mathrm{d}\gamma^{\prime\prime}/\mathrm{d}\omega]|_{\omega\rightarrow 0}$ (the sloped dash-dotted line) intercepts the horizontal lines $\gamma(0)$ and $\gamma_{a}$. The difference of the abscissas is just $\Delta(\omega\tau)=1.$ This is the tool to determine the time constant $\tau$ according Eq. (130) and Eq. (131).
The real part $\gamma^{\prime}(\omega)$ monotonously increases from the thermodynamically calculated $\gamma_{0}$ to the monoatomic value $\gamma_{a}$. The imaginary part $\gamma^{\prime\prime}(\omega\tau)$ is negligible at very low and very high frequencies. It has a maximum marked by dot at frequency for which $\omega\tau=1$. Dots at the $\gamma^{\prime}(\mathcal{W})$ dependence and the tangent at the $\gamma^{\prime\prime}(\mathcal{W})$ illustrate the determination of time-constant $\tau$ according to Eq. (131).
Perhaps in astrophysical problems the bulk viscosity could be the most important for calculation of acoustic heating by the space damping rate $k^{\prime\prime}(f)$ as a function of the frequency. This dependence is represented in logarithmic scale in Fig. 8.
Figure 8: Imaginary part of the wave-vector determining wave amplitude damping $k^{\prime\prime}(f=\omega/2\pi)$ in homogeneous fluid $\propto\mathrm{e}^{-k^{\prime\prime}x}$. At low frequencies space damping rate $k^{\prime\prime}\propto\omega^{2}$ according Eq. (132). At high frequencies $\omega\tau\gg 1$ damping rate reaches high-frequency asymptotic $k_{\infty}^{\prime\prime}$ given by Eq. (133).
For extremely low frequencies we have power law $k^{\prime\prime}\propto f^{2}$ according to Eq. (132). For higher frequencies, the damping rate is dispersion-less $k^{\prime\prime}\approx\mathrm{const}$ according to Eq. (133). This is an important simplification: in the frequency interval $\omega\in(1/\tau,\,\sqrt{\mathrm{P}_{\zeta/\eta}}/\tau)$ the spectral density of the heating power is proportional to the spectral density of the acoustic waves; see the explanation below.
Finally we have to compare the high frequency behavior of the bulk viscosity $\zeta^{\prime}(\omega)\approx\zeta_{0}/(\omega\tau)^{2}$ with the shear one which has no frequency dependent dispersion. They are equal at the critical frequency $f_{c}=\omega_{c}/2\pi$ for which
$\zeta^{\prime}(\omega_{c})=\eta,\quad\omega_{c}=\frac{1}{\tau}\sqrt{\mathrm{P}_{\zeta/\eta}},\quad\mathrm{P}_{\zeta/\eta}\equiv\frac{\zeta_{0}}{\eta}\gg 1.$ (137)
In order the short wavelength approximation to be applicable and the plasma to be considered as a homogeneous medium, the wavelength must be smaller than the typical length $h$ of the system
$\lambda_{c}\equiv c_{s}/f_{c}\ll h.$ (138)
For frequencies smaller than $f_{c}$, the bulk viscosity dominates in the wave damping and acoustic heating. In our illustrative example from the AL08 profile
$c_{s}\approx 9.24~\mathrm{km/s},\quad f_{c}\approx 35~\mathrm{mHz},\quad\lambda_{c}\approx 264~\mathrm{km}$ (139)
at $h=1894$ km. Our qualitative conclusion is that the bulk viscosity indispensably has to be included in the models for acoustic heating of the solar chromosphere [1, 2, 3, 4, 5] and models involving low temperature plasmas [24] if sound waves have significant spectral density for frequencies $\ll f_{c}$.
VI Hydrogen-helium (alkali-noble) cocktail analytical solution for the thermodynamics and kinetics
In order to illustrate the work of the general scheme, here we analyze the important special case of a cold cocktail of partially ionized hydrogen and neutral helium atoms for the solar chromosphere, i.e.
$\epsilon\equiv\epsilon_{\mathrm{H}}\equiv\epsilon_{1},\qquad\epsilon_{\mathrm{He}}\equiv\epsilon_{2}=0,$ (140)
$\overline{a}_{\mathrm{H}}=1,\qquad\overline{a}_{\mathrm{He}}=0.1\,$, (141)
$\rho=M^{*}n_{\rho},\qquad M^{*}=M+\overline{a}_{\mathrm{He}}M_{\mathrm{He}},$ (142)
$1/\tau_{\mathrm{{}_{H}}}=1/\tau_{1}=\beta(T)n_{e}=\beta n_{\rho}\overline{\alpha},$ (143)
$\overline{\mathcal{N}}_{\mathrm{tot}}=1+\overline{\alpha}+\overline{a}_{\mathrm{He}},$ (144)
$\overline{\alpha}(T)\equiv\overline{\alpha}_{\mathrm{H}}=\dfrac{2}{1+\sqrt{4\dfrac{\overline{n}_{\rho}}{n_{q}}+1}},\qquad\overline{\alpha}_{\mathrm{He}}=0.$ (145)
And for the pressure in the approximation of ideal gas we have
$p=n_{\mathrm{tot}}T=\mathcal{N}_{\mathrm{tot}}n_{\rho}T=(1+\alpha+\overline{a}_{\mathrm{He}})n_{\rho}T.$ (146)
The Saha equation can be written as
$\alpha(\rho,T)=f(\nu)\equiv\alpha_{\mathrm{H}}\equiv\dfrac{2}{1+\sqrt{1+4\nu}},$ (147)
$\nu\equiv\dfrac{n_{\rho}}{n_{\mathrm{{}_{S}}}}=\frac{1-\alpha}{\alpha^{2}},\qquad\dfrac{1}{\sqrt{1+4\nu}}=\dfrac{\alpha}{2-\alpha}$ (148)
$n_{\mathrm{{}_{S}}}(T)\equiv n_{q}\mathrm{e}^{-\iota},\quad n_{q}(T)=\left(\dfrac{mT}{2\pi\hbar^{2}}\right)^{\!\!3/2},\quad\iota\equiv\dfrac{I}{T},$
where $m$ is the electron mass and $n_{\rho}=\rho/M^{*}$, Its solution for the hydrogen degree of ionization $\overline{\alpha}(T,\overline{n}_{\rho})$ is given in Ref. [21]. Actually, this is an applicable approximation for the solar chromosphere but the derived formulas are applicable for any alkali-noble cocktail, Na-Ne plasma or arbitrary cold pure element plasma.
VI.1 Thermodynamics of the cold binary plasma
For the enthalpy $w$ and internal energy $\varepsilon$ per unit mass we have [25, 13, Eqs. (11-12), Eq. (6)]
$w(\rho,T)$ $=\dfrac{1}{M^{*}}\left(c_{p}\mathcal{N}_{\mathrm{tot}}T+I\alpha\right),\qquad c_{p}=\dfrac{5}{2},$ (149)
$\varepsilon(\rho,T)$ $=\dfrac{1}{M^{*}}\left(c_{v}\mathcal{N}_{\mathrm{tot}}T+I\alpha\right),\qquad c_{v}=\dfrac{3}{2}.$ (150)
Here we wish to emphasize that magnetic field has zero influence on the thermodynamics of the solar atmosphere which in a classical gas. This statement is known as Bohr–van Leeuwen [26] theorem. For the derivatives we have
$f^{\prime}(\nu)$ $=\mathrm{d}_{\nu}f(\nu)=-\dfrac{4}{\left(1+\sqrt{1+4\nu}\right)^{\!2}\sqrt{1+4\nu}}$
$=\dfrac{-f^{2}(\nu)}{\sqrt{1+4\nu}}=\dfrac{-f^{3}}{2-f}=-\dfrac{\alpha^{3}}{2-\alpha}=\dfrac{\mathrm{d}\alpha}{\mathrm{d}\nu},$ (151)
$\nu f^{\prime}(\nu)$ $=-\dfrac{(1-\alpha)\alpha}{2-\alpha}=-\mathcal{D},\quad\mathcal{D}\equiv\dfrac{(1-\alpha)\alpha}{2-\alpha},$ (152)
$\left(\dfrac{\partial\alpha}{\partial T}\right)_{\!\!\rho}$ $=-\dfrac{(c_{v}+\iota)}{T}\nu f^{\prime}(\nu)=\dfrac{(c_{v}+\iota)}{T}\dfrac{(1-\alpha)\alpha}{2-\alpha}$
$=\dfrac{(c_{v}+\iota)}{T}\mathcal{D},$ (153)
$\left(\dfrac{\partial\alpha}{\partial\rho}\right)_{\!T}$ $=\nu f^{\prime}(\nu)/(M^{*}n_{\mathrm{{}_{S}}})=-\dfrac{1}{M^{*}n_{\mathrm{{}_{S}}}}\dfrac{(1-\alpha)\alpha}{2-\alpha}$
$=-\dfrac{1}{M^{*}n_{\mathrm{{}_{S}}}}\mathcal{D}.$ (154)
In such a way we obtain for the heat capacity the explicit expressions
$\mathcal{C}_{v}$ $\equiv\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{\!\!\rho}=\dfrac{1}{M^{*}}\left[c_{v}\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)^{2}\mathcal{D}\right]$ (155)
$=\dfrac{1}{M^{*}}\left[c_{v}(1+\alpha+\overline{a}_{\mathrm{He}})+(c_{v}+\iota)^{2}\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]$ (156)
$=T\left(\dfrac{\partial s}{\partial T}\right)_{\!\!\!\rho},$
where $s$ is the entropy per unit mass.
For derivatives of the pressure Eq. (146) we have
$\left(\dfrac{\partial p}{\partial T}\right)_{\!\!\rho}$ $=\left[\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)\mathcal{D}\right]n_{\rho}$ (157)
$=\left[(1+\alpha+\overline{a}_{\mathrm{He}})+(c_{v}+\iota)\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]n_{\rho},$
$\left(\dfrac{\partial p}{\partial\rho}\right)_{\!\!T}$ $=\dfrac{T}{M^{*}}\left[\mathcal{N}_{\mathrm{tot}}-\mathcal{D}\right]$ (158)
$=\dfrac{T}{M^{*}}\left[(1+\alpha+\overline{a}_{\mathrm{He}})-\dfrac{(1-\alpha)\alpha}{2-\alpha}\right].$
Heat capacity per constant volume and unit mass $\mathcal{C}_{v}$ can be obtained by the general formula [15, Eq. (16.10)] in which volume is for unit mass $\mathcal{V}\equiv 1/\rho$
$\Delta\mathcal{C}$ $\equiv\mathcal{C}_{p}-\mathcal{C}_{v}=-T\dfrac{(\partial p/\partial T)_{\mathcal{V}}^{2}}{(\partial p/\partial\mathcal{V})_{T}}=\dfrac{T}{\rho^{2}}\dfrac{(\partial p/\partial T)_{\rho}^{2}}{(\partial p/\partial\rho)_{T}}$
$=\dfrac{\left[\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)\mathcal{D}\right]^{2}}{M^{*}\left[\mathcal{N}_{\mathrm{tot}}-\mathcal{D}\right]}$ (159)
$=\dfrac{\left[(1+\alpha+\overline{a}_{\mathrm{He}})+(c_{v}+\iota)\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]^{2}}{M^{*}\left[(1+\alpha+\overline{a}_{\mathrm{He}})-\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]},$
and
$\mathcal{C}_{p}=$ $\dfrac{\left[c_{v}\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)^{2}\mathcal{D}\right]}{M^{*}}+\dfrac{\left[\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)\mathcal{D}\right]^{2}}{M^{*}\left[\mathcal{N}_{\mathrm{tot}}-\mathcal{D}\right]}$
$=$ $\dfrac{\left[c_{p}\mathcal{N}_{\mathrm{tot}}+(c_{v}c_{p}+2c_{p}\iota+\iota^{2})\mathcal{D}\right]\mathcal{N}_{\mathrm{tot}}}{(\mathcal{N}_{\mathrm{tot}}-\mathcal{D})M^{*}}$ (160)
$=$ $\dfrac{1}{M^{*}}\left[c_{v}(1+\alpha+\overline{a}_{\mathrm{He}})+(c_{v}+\iota)^{2}\frac{(1-\alpha)\alpha}{2-\alpha}\right]$
$+\dfrac{\left[(1+\alpha+\overline{a}_{\mathrm{He}})+(c_{v}+\iota)\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]^{2}}{M^{*}\left[(1+\alpha+\overline{a}_{\mathrm{He}})-\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]}$ (161)
$=T\left(\dfrac{\partial s}{\partial T}\right)_{\!\!p}.$ (162)
These results generalize the previously obtained heat capacities for pure hydrogen [25].
Analogously for the derivatives of enthalpy
$\left(\dfrac{\partial w}{\partial T}\right)_{\!\!\rho}$ $=\dfrac{1}{M^{*}}\left[c_{p}\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)(c_{p}+\iota)\mathcal{D}\right]$ (163)
$=\left[(1+\alpha+\overline{a}_{\mathrm{He}})c_{p}+(c_{v}+\iota)(c_{p}+\iota)\dfrac{(1-\alpha)\alpha}{2-\alpha}\right],$
$\left(\dfrac{\partial w}{\partial\rho}\right)_{\!\!T}$ $=-\dfrac{T}{M^{*}}(c_{p}+\iota)\dfrac{\mathcal{D}}{M^{*}n_{\rho}}$ (164)
$=-\dfrac{T}{M^{*}}\dfrac{1}{M^{*}n_{\rho}}\left[(c_{p}+\iota)\dfrac{(1-\alpha)\alpha}{2-\alpha}\right].$
Then the Jacobian [13, Eq. (9)] with dimension of reciprocal mass
$\mathcal{J}\equiv\frac{\partial(w,p)}{\partial(T,\rho)}=\left(\frac{\partial w}{\partial T}\right)_{\!\!\rho}\left(\frac{\partial p}{\partial\rho}\right)_{\!\!T}-\left(\frac{\partial w}{\partial\rho}\right)_{\!\!T}\left(\frac{\partial p}{\partial T}\right)_{\!\!\rho}$ (165)
after some algebra reads
$\mathcal{J}=\frac{T\mathcal{N}_{\mathrm{tot}}}{(M^{*})^{2}}\left\{c_{p}\mathcal{N}_{\mathrm{tot}}+\left[(c_{v}+\iota)(c_{p}+\iota)+\iota\right]\mathcal{D}\right\}=\frac{(1\!+\!\alpha\!+\!\overline{a}_{\mathrm{He}})T}{(M^{*})^{2}}\left[c_{p}(1\!+\!\alpha\!+\!\overline{a}_{\mathrm{He}})+(c_{v}c_{p}+2c_{p}\iota+\iota^{2})\frac{(1-\alpha)\alpha}{2-\alpha}\right].$ (166)
In such a way we obtained that for the 2-component cocktail
$\mathcal{J}=\frac{T}{M^{*}}(\mathcal{N}_{\mathrm{tot}}-\mathcal{D})\mathcal{C}_{p}.$ (167)
Representing the sound velocity at evanescent frequency when ionization degree follows it equilibrium values determined by Saha equation we have
$c_{0}^{2}\equiv\left(\frac{\partial p}{\partial\rho}\right)_{\!\!s}=\gamma_{0}\frac{p}{\rho},$ (168)
The general result [13, Eq. (9)]
$\gamma_{0}=\frac{\rho}{p}\cdot\frac{\mathcal{J}}{\mathcal{C}_{p}}$ (169)
after substitution $\mathcal{J}$ from Eq. (167) gives
$\gamma_{0}$ $=$ $\left(1-\dfrac{\mathcal{D}}{\mathcal{N}_{\mathrm{tot}}}\right)\dfrac{\mathcal{C}_{p}}{\mathcal{C}_{v}}$
$=$ $\left[1-\dfrac{(1-\alpha)\alpha}{(2-\alpha)(1+\alpha+\overline{a}_{\mathrm{He}})}\right]\dfrac{\mathcal{C}_{p}}{\mathcal{C}_{v}}.$
This illustrates how the ionization recombination processes break the general relations derived for constant chemical compounds. For solar plasma it is not so strongly expressed but for interstellar plasma where $\iota=I/T\gg 1$ it is possible $\iota^{2}\mathcal{D}$ to dominate, hence
$M^{*}\mathcal{C}_{v}$ $\approx$ $\iota^{2}\mathcal{D}\gg 1,$ (171)
$M^{*}\mathcal{C}_{p}$ $\approx$ $\dfrac{\iota^{2}\mathcal{D}}{1-\mathcal{D}/\mathcal{N}_{\mathrm{tot}}}\gg 1,$ (172)
confer Refs. [25, 21, Fig. 3 and Fig. 5]. In the opposite case of negligible influence of the ionization processes, $\iota\ll 1$ we have the test for the programming of the heat capacities
$M^{*}\mathcal{C}_{v}\approx c_{v}\mathcal{N}_{\mathrm{tot}},\;M^{*}\mathcal{C}_{p}\approx c_{p}\mathcal{N}_{\mathrm{tot}},\;M^{*}(\mathcal{C}_{p}-\mathcal{C}_{v})=\mathcal{N}_{\mathrm{tot}}.$ (173)
The general formula for pure hydrogen plasma was solved in the early stages of the present study [21, Eq. (96)] and [25, Eq. (17)]
$c_{0}^{2}=\frac{2c_{p}+(1-\alpha)\alpha(c_{p}+\iota)^{2}}{2c_{v}+(1-\alpha)\alpha\left[(c_{v}+\iota)^{2}+c_{v}\right]}\frac{\bar{p}}{\bar{\rho}},$ (174)
where $\iota\equiv I/T$, $I=13.6$ eV is the hydrogen ionization potential, $\alpha\equiv\overline{\alpha}$ is the equilibrium ionization degree
$\alpha\equiv\frac{n_{e}}{n_{p}+n_{0}}=\frac{n_{e}}{n_{\rho}}\in(0,1)\,$, (175)
$n_{e}=n_{p}=\alpha n_{\rho},\quad n_{0}=(1-\alpha)\,n_{\rho},\quad$ (176)
$n_{\mathrm{tot}}=n_{e}+n_{p}+n_{0}=(1+\alpha)\,n_{\rho},$ (177)
where $n_{e}$, $n_{p}$ and $n_{0}$ are the concentrations of electrons, protons and neutral atoms respectively, and we denote the sum of the concentrations of the protons and the neutrals by $n_{\rho}=n_{0}+n_{p}$. This formula can be generalized for arbitrary low temperature cocktail of alkaline metal and noble gas [13, Eq. (B1)], (for example Na-Ne plasma)
$\gamma_{0}=\frac{c_{p}[2+(2-\alpha)\overline{a}_{\mathrm{nbl}}]+(1-\alpha)\alpha(c_{p}+\iota)^{2}}{c_{v}[2+(2\!-\!\alpha)\overline{a}_{\mathrm{nbl}}]+(1\!-\!\alpha)\alpha\left[(c_{v}+\iota)^{2}+c_{v}\right]},$ (178)
where
$\overline{a}_{\mathrm{nbl}}=\frac{n_{\mathrm{nbl}}}{n_{\mathrm{alk}}},\quad\alpha\equiv\overline{\alpha}=\overline{r}_{1,\,\mathrm{alk}},\qquad I=I_{1,\,\mathrm{alk}}.$ (179)
VI.2 Kinetics of the cold binary plasma
In this special case the general Eq. (76) is reduced to a simple linear equation for $\epsilon$ having the solution
$\epsilon=\dfrac{\dfrac{\iota}{c_{v}}\epsilon_{\rho}}{\dfrac{1-\mathrm{i}\omega\tau_{1}\overline{\alpha}}{1-\overline{\alpha}}+\dfrac{\overline{\alpha}}{\overline{\mathcal{N}}_{\mathrm{tot}}}\left[\left(\dfrac{\iota}{c_{v}}+2\right)\!\iota+c_{v}+\dfrac{1}{\overline{c}_{e}}\right]}.$ (180)
The substitution of this solution in the final expression for the generalized polytropic index Eq. (96) gives
$\hat{\gamma}(\omega)$ $=\frac{c_{p}}{c_{v}}-\dfrac{\left(\dfrac{\iota}{c_{v}}\right)^{\!2}(1-\overline{\alpha})\,\overline{\alpha}}{(1-\mathrm{i}\omega\tau_{\mathrm{{}_{H}}}\overline{\alpha})(1+\overline{\alpha}+\overline{a}_{\mathrm{He}})+\left[\left(\dfrac{\iota}{c_{v}}+2\right)\!\iota+c_{v}+\dfrac{1+\overline{\alpha}+\overline{a}_{\mathrm{He}}}{\overline{\alpha}}\right](1-\overline{\alpha})\,\overline{\alpha}}$ (181)
$=\gamma_{a}-\dfrac{(1-\overline{\alpha})\,\overline{\alpha}\,\iota^{2}/c_{v}}{(1-\mathrm{i}\omega\tau_{\mathrm{{}_{H}}}\overline{\alpha})c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}+\left[(c_{v}+\iota)^{2}\overline{\alpha}+c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}\right](1-\overline{\alpha})}$ (182)
$=\frac{c_{p}}{c_{v}}-\dfrac{\left(\dfrac{\iota}{c_{v}}\right)^{\!2}(1-\overline{\alpha})\,\overline{\alpha}}{(1-\mathrm{i}\omega\tau_{\mathrm{{}_{H}}}\overline{\alpha})(1+\overline{\alpha}+\overline{a}_{\mathrm{He}})+\left[\dfrac{(c_{v}+\iota)^{2}}{c_{v}}+\dfrac{1+\overline{\alpha}+\overline{a}_{\mathrm{He}}}{\overline{\alpha}}\right](1-\overline{\alpha})\,\overline{\alpha}}.$ (183)
This solution for the generalized polytropic index for the alkali-noble cocktail Eq. (182) can be rewritten as
$\hat{\gamma}(\omega)=\gamma_{\infty}-\frac{\mathcal{B}}{\mathcal{A}-\mathrm{i}\omega\mathcal{T}},$ (184)
$\mathcal{A}\equiv(2-\overline{\alpha})c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}+(c_{v}+\iota)^{2}(1-\overline{\alpha})\overline{\alpha}$ (185)
$\mathcal{B}=(1-\overline{\alpha})\,\overline{\alpha}\,\iota^{2}/c_{v}$ (186)
$\mathcal{T}\equiv\tau_{\mathrm{{}_{H}}}c_{v}\overline{\alpha}\overline{\mathcal{N}}_{\mathrm{tot}},$ (187)
and in these notations
$\gamma^{\prime\prime}=-\frac{\mathcal{B}\,\omega\mathcal{T}}{\mathcal{A}^{2}+(\omega\mathcal{T})^{2}},\qquad-\left.\frac{\gamma^{\prime\prime}(\omega)}{\omega}\right|_{\omega\rightarrow 0}=\frac{\mathcal{B}\mathcal{T}}{\mathcal{A}^{2}}.$ (188)
It is instructive to note that
$\mathcal{\overline{N}}_{\mathrm{tot}}$ $=1+\overline{a}_{\mathrm{He}}+\overline{\alpha}\equiv\overline{a}_{\mathrm{H}}+\overline{a}_{\mathrm{He}}+\overline{a}_{e}$
$=\sum_{a}^{a_{\mathrm{max}}}\overline{a}_{a}+\overline{a}_{e},\quad\overline{a}_{\mathrm{H}}=1,\quad\overline{a}_{e}=\overline{\alpha},$ (189)
clearly giving the recipe to include more noble gases in the cocktail. Additionally from Eq. (184) we have
$\gamma_{\infty}^{\prime}-\gamma_{0}^{\prime}=\gamma_{a}-\gamma(\omega=0)=\mathcal{B}/\mathcal{A}$ (190)
and the substitution in Eq. (130) gives for the Mandelstam-Leontovich time-constant
$\tau$ $=\frac{\mathcal{T}}{\mathcal{A}}=\frac{\overline{\alpha}c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}}{\mathcal{A}}\tau_{\mathrm{{}_{H}}}$ (191)
$=\frac{c_{v}\overline{\alpha}\overline{\mathcal{N}}_{\mathrm{tot}}}{(2-\overline{\alpha})c_{v}\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)^{2}(1-\overline{\alpha})\overline{\alpha}}\tau_{\mathrm{{}_{H}}}$ (192)
$=\frac{(1\!+\!\overline{a}_{\mathrm{He}}\!+\!\overline{\alpha})\overline{\alpha}c_{v}}{(2\!-\!\overline{\alpha})(1\!+\!\overline{a}_{\mathrm{He}}\!+\!\overline{\alpha})c_{v}+(c_{v}\!+\!\iota)^{2}(1\!-\!\overline{\alpha})\overline{\alpha}}\tau_{\mathrm{{}_{H}}}.$ (193)
In agreement with the result for pure hydrogen plasma  [21, Eq. (113)] derived in a different way. This time constant substituted in Eq. (128) gives, cf. [21, Eq. 108]
$\zeta_{0}$ $=\frac{\overline{p}\mathcal{B}\mathcal{T}}{\mathcal{A}^{2}}=\frac{\overline{p}\mathcal{B}\overline{\alpha}c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}\tau_{\mathrm{{}_{H}}}}{\mathcal{A}^{2}}$ (194)
$=\frac{(1-\overline{\alpha})\,\overline{\alpha}\,\iota^{2}(1+\overline{a}_{\mathrm{He}}+\overline{\alpha})\overline{p}\tau_{\mathrm{{}_{H}}}}{[(2-\overline{\alpha})(1+\overline{a}_{\mathrm{He}}+\overline{\alpha})c_{v}+(c_{v}+\iota)^{2}(1-\overline{\alpha})\overline{\alpha}]^{2}}.$
For the frequency dependent sound velocity $c_{\omega}$ the ML theory gives the real part of a complex one [16, Eq. (81.8)]
$c_{\omega}\approx\Re(\hat{c}_{\mathrm{phase}})=\frac{\omega}{\Re(\hat{k})}=\Re\left(\sqrt{\frac{c_{0}^{2}-\mathrm{i}\omega c_{\infty}^{2}}{1-\mathrm{i}\omega\tau}}\right),$ (195)
$\hat{k}=\omega\sqrt{\frac{1-\mathrm{i}\omega\tau}{c_{0}^{2}-\mathrm{i}\omega c_{\infty}^{2}}}.$ (196)
The purpose of the present study is to represent how the time-constant $\tau$ can be obtained by solving the kinetic equation with ab initio calculated ionization cross-section and sound velocity $c_{0}$ is derived from the thermodynamic approach applied to mono-atomic plasma cocktail.
Finally the important for us high-frequency dispersion-less damping according Eq. (133)
$2k_{\infty}^{\prime\prime}$ $\equiv 2k^{\prime\prime}(\omega\rightarrow\infty)=\frac{1}{\tau c_{\infty}}\frac{\gamma_{a}-\gamma_{0}}{\gamma_{a}}=\frac{1}{\tau c_{\infty}}\frac{\mathcal{B}}{\gamma_{a}\mathcal{A}}$ (197)
$=\frac{\beta(T)n_{\rho}\iota^{2}(1-\overline{\alpha})\overline{\alpha}}{(1+\overline{\alpha}+\overline{a}_{\mathrm{He}})c_{\infty}c_{p}c_{v}}$ (198)
$=\frac{(\gamma_{a}-1)^{2}(1-\overline{\alpha})\overline{\alpha}\iota^{2}}{\gamma_{a}\overline{\mathcal{N}}_{\mathrm{tot}}}\frac{1}{\tau_{\alpha}c_{\infty}},\quad\tau_{\alpha}=1/\beta n_{\rho}$ (199)
$\rightarrow\sum_{a}\frac{I_{a}^{2}}{c_{p}c_{v}T^{2}}\frac{\langle\sigma_{a}v\rangle}{c_{\infty}}\frac{\overline{n}_{e}\overline{n}_{0,a}}{\overline{n}_{\mathrm{tot}}}\equiv\frac{1}{l},$ (200)
$c_{\infty}$ $=\sqrt{\frac{\gamma_{a}\overline{p}}{\overline{\rho}}},\quad\gamma_{a}=\gamma_{\infty}=\frac{5}{3},\quad c_{p}=\gamma_{a}c_{v},\quad c_{v}=\frac{3}{2}.$
See also the result for pure hydrogen plasma [21, Eq. (112)] The last expression akin of Clausius concept for the mean free path $l\sigma n\sim 1$. The reciprocal mean free path $1/l$ is additive and sum on all kind scattering centers, in our case on all kind neutral atoms.
In the general case we have $\hat{k}(\omega)$ expressed by Eq. (101) with $\hat{\gamma}(\omega)$ from Eq. (184). In such a way our general conclusion is that in alkaline-noble gas approximation we have analytical expression for $\zeta_{0}$ and $\tau$ and Mandelstam-Leontovich formula for $\zeta(\omega)$ gives the exact frequency dependence.
VI.3 Special cases analysis
  • For completely ionized hydrogen plasma $\overline{\alpha}=1$ we have $\hat{\gamma}(\omega)=\gamma_{a}$.
  • For atomic hydrogen $\overline{\alpha}=0$ also because for monoatomic gases the bulk viscosity is zero [27].
  • For very high frequencies $\omega\rightarrow\infty$, $\hat{\gamma}(\omega)=\gamma_{a}$ and $\zeta^{\prime}=0$ again because the influence of the ionization-recombination processes in this $\omega\tau\gg 1$ case is negligible.
  • In the opposite case of evanescent frequencies at formally $\omega=0$, we reproduce after some algebra $\hat{\gamma}(\omega=0)=\gamma_{0}$ the thermodynamic result Eq. (178[13, Eq. (B1)], with $\overline{a}_{\mathrm{nbl}}\equiv\overline{a}_{\mathrm{He}}$ and $\alpha\equiv\overline{\alpha}$.
  • For interstellar plasma it is possible $T\ll I$ at significant $(1-\overline{\alpha})\overline{\alpha}$, formally at $(1-\overline{\alpha})\overline{\alpha}\iota^{2}\gg 1$ we obtain $\gamma_{0}\approx\gamma_{a}-1/c_{v}=1$, $c_{0}=c_{\mathrm{N}}$ and $c_{0}^{2}=\overline{p}/\overline{\rho}$. This formula suggested initially by Newton is an important test for the general program for numerical calculation of bulk viscosity of cold plasma.
  • As a function of the frequency $\omega$ Eq. (181) gives a simple algebraic expression linear function divided by another linear function and taking into account the limit cases at low and high frequencies, we conclude that in this singly ionized atomic approximation we arrive at Mandelstam-Leontovich formula Eq. (103) which we have derived in the present study by the microscopic kinetic approach.
  • Finally, knowing the general relation between the bulk viscosity and generalized polytropic index Eq. (123), we can determine the relaxation time $\tau$ of the ML theory. For example, comparing the high frequency behavior Eq. (136) with the low frequency approximation Eq. (130), we again arrive at Eq. (193).
  • Having a reliable approximation for low temperature hydrogen-helium cocktail, we can address some practical for plasma physics problems as the long standing problem of heating of solar chromosphere, for example.
VI.4 Solar chromospheric profiles of the main notions and damping of magneto-hydrodynamic waves
Using the semi-empirical height dependent profiles of the temperature $T^{\prime}$ and mass density $\rho$ [6, Model C7] (AL08), the main introduced notions are calculated and analyzed in this subsection.
First, the height profile of the function $\mathcal{D}(h)$ is depicted in Fig. 9 below the profile of $\iota(h)$.
Figure 9: The height dependence of dimensionless functions $\lg{[\iota(h)\equiv I/T(h)]}$ (solid line) and $\lg\mathcal{D}(h)$ (dashed line) Eq. (152) which is via Ref. [6, Model C7] in the ordinate for heights $h\equiv x<2.5$ Mm in the abscissa. The function $\mathcal{D}$ is the main ingredient of all analytical results in which ionization-recombination processes are relevant. For comparison in the same logarithmic ordinate is given the profile of the dimensionless variable $\iota(h)\gg 1$. The abrupt change of both variables is physically in the solar transition region and it is beyond the scope of the present study.
Both these notions are central ingredients in our analytical results. As $\mathcal{D}\rightarrow 0$ in the solar transition region, $\zeta\rightarrow 0$ and $\iota<1$ in the coronal conditions beyond the current study.
The comparison between different sound waves damping mechanisms represented in Ref. [28, Fig. 1] for the solar atmosphere via AL08 reveals that the bulk viscosity is many orders of magnitude larger than the shear one in Fig. 10 where the height dependence of bulk viscosity Prandtl number
$\mathrm{P}_{\zeta/\eta}\equiv\frac{\zeta_{0}}{\eta}\gg 1.$ (201)
is shown.
Figure 10: The height $h$ profile of bulk viscosity Prandtl number Eq. (201) $\mathrm{P}_{\zeta/\eta}\equiv\zeta_{0}/\eta$ via Avrett and Loeser [6, Model C7]. One of the purposes of the present study is to emphasize that for heights $h\equiv x<2.1$ Mm from the solar photospere the bulk viscosity indispensable must be included in the considerations of acoustic wave heating of chromosphere. For some heights viscous Prandtl number $\mathrm{P}_{\zeta/\eta}$ can reach $10^{11}$; neglecting of $\zeta$ on the background of $\eta$ in this case is equivalent of neglecting of a giraffe on the background of a hydrogen atom.
The collisions of protons with protons are more intensive than collisions with neutral atoms and for the shear viscosity we use the result for completely ionized hydrogen plasma [27, Eq. (43.9)]
$\eta\approx 0.4\dfrac{M^{1/2}T^{5/2}}{e^{4}\Lambda},\qquad\Lambda=\ln\dfrac{Tr_{\mathrm{{}_{D}}}}{e^{2}},$ (202)
$e^{2}=\dfrac{q_{e}^{2}}{4\pi\varepsilon_{0}},\qquad\dfrac{1}{r_{\mathrm{{}_{D}}}^{2}}=2\,\dfrac{4\pi e^{2}}{T}n_{\rho}\alpha,$ (203)
where $q_{e}$ is the electron charge, and $r_{\mathrm{{}_{D}}}$ is the Debye radius [15, Eq. (78.8)]. Analogously, the electrons are scattered mainly by protons and [27, Eq. (43.10)]
$\varkappa=\dfrac{T^{5/2}}{e^{4}m^{1/2}\Lambda}.$ (204)
Both kinetic coefficients $\eta$ and $\varkappa$ have weak density dependence only through the Coulomb logarithm $\Lambda$. The total space damping rate of sound (acoustic) waves [16, Eq. (79.6)]
$k^{\prime\prime}(\omega)=\frac{\omega^{2}}{2\rho c^{3}}\left[\left(\frac{4}{3}\eta+\zeta^{\prime}(\omega)\right)+\left(\frac{1}{\mathcal{C}_{v}}-\frac{1}{\mathcal{C}_{p}}\right)\varkappa\right]$ (205)
depends on shear $\eta$ and frequency dependent bulk viscosity $\zeta^{\prime}(\omega)$, thermal conductivity and heat capacities per unit mass at constant volume $\mathcal{C}_{v}$ and pressure $\mathcal{C}_{p}$. For the coefficient of heat conductivity damping in Eq. (205) we have an explicit and computable expression
$\frac{1}{\mathcal{C}_{v}}-\frac{1}{\mathcal{C}_{p}}=\frac{\Delta\mathcal{C}}{\mathcal{C}_{v}\mathcal{C}_{p}}.$ (206)
The same approach allows the incorporation of magneto-hydrodynamic (MHD) or Alfvén waves (AW) [29, 30] with vertical magnetic field in the same scheme. For the AW with the dispersion [31, Eq. (69.8)] supposing that magnetic field is vertical and parallel to the wave-vector
$\omega_{\mathrm{{}_{AW}}}=V_{\mathrm{A}}k_{\mathrm{{}_{AW}}}^{\prime},\qquad V_{\mathrm{A}}=\frac{B}{\sqrt{\mu_{0}\rho}},\qquad\mu_{0}=4\pi$ (207)
the damping rate [31, Problem to Sec. 69]
$2k_{\mathrm{{}_{AW}}}^{\prime\prime}$ $=(\nu_{k}+\nu_{m})\,\frac{\omega^{2}}{V_{\mathrm{A}}^{3}},\qquad\nu_{k}\equiv\frac{\eta}{\rho},$ (208)
$\nu_{m}$ $\equiv\varepsilon_{0}c^{2}\varrho_{{}_{\Omega}}=\frac{c^{2}}{4\pi}\frac{e^{2}m^{1/2}\Lambda}{0.6\,T^{3/2}},$ (209)
where $\varepsilon_{0}=1/4\pi$, $c$ is the light velocity, and $\varrho_{{}_{\Omega}}$ is the Ohmic resistivity [32, Eq. (43.8-10)]. The comparison of Alfvén wave (AW) damping Eq. (208) with acoustic wave damping Eq. (205) shows that in some circumstances partial ionized plasma could operate as polarizer absorbing longitudinal acoustic waves and transmitting transversal AW. Both the formulae Eq. (205) and Eq. (208) for the wave damping can be derived as ratio of heat production (energy dissipation) $Q_{\mathrm{wave}}$ divided by the energy flux of a plane wave with fixed frequency $\omega$
$2k^{\prime\prime}=\mathcal{Q}_{\mathrm{wave}}/q_{\mathrm{wave}}.$ (210)
Even strong magnetic field when parallel to wave-vector has no influence on the dissipative kinetic coefficients describing wave damping.
For high enough frequencies the viscosity terms are equalized $\zeta_{0}/(\omega_{\eta}\tau)^{2}=\eta$ and above $\omega_{\eta}\equiv 2\pi f_{\eta}$ the frequency independent shear viscosity is larger. In Fig. 11 the height $h$ profile of the frequency interval $(f_{\mathrm{ML}},\,f_{\eta})$ is depicted where the damping rate $k^{\prime\prime}$ is almost frequency independent.
Figure 11: Height $h$ profiles of the frequencies $f_{\mathrm{ML}}\equiv 1/2\pi\tau\ll f\ll f_{\eta}\equiv f_{\mathrm{ML}}\sqrt{\mathrm{P}_{\zeta/\eta}}$ again via Avrett and Loeser [6, Model C7]. The damping rate $k_{\infty}$ is frequency independent according to Eq. (200) and determined by the bulk viscosity $\zeta$ and volume heating.
This dramatically simplifies the theoretical analysis; no need to study the frequency dependence of the spectral density $\varrho(f)$ of the acoustic waves
$q=\int\limits_{0}^{\infty}\varrho(f)\mathrm{d}f.$ (211)
At the height of the solar transition region $f_{\mathrm{ML}}$ and $f_{\eta}$ exchange their places as $\alpha\approx 1$ leading to $\mathcal{D}\rightarrow 0$ in Fig. 9, meaning that the bulk viscosity heating is abruptly terminated there.
As a concluding remark of this section, it is good to repeat that formulae derived for the hydrogen-helium cocktail of the solar plasma are applicable for laboratory alkaline-noble gas cocktails.
VI.5 Entropy production
Here we summarize the fundament of the kinetics on which is build the theory of bulk viscosity and wave damping in general. The main detail of the kinetics is the entropy production and for hydrodynamics the heating per unit volume. Let us recall [16, Eq. (49.6)] in the beginning the influence of the shear viscosity and corresponding incompressible motion of the fluid with velocity vector $\mathbf{v}(t.\mathbf{r})$
$Q_{\eta}=\frac{\eta}{2}\sum_{i,\,j}\left(\frac{\partial v_{i}}{\partial x_{k}}+\frac{\partial v_{k}}{\partial x_{i}}-\frac{2}{3}\delta_{i,k}\sum_{l}\frac{\partial v_{l}}{\partial x_{l}}\right)^{\!\!\!2}.$ (212)
Analogously for the bulk viscosity heating low frequency formula
$Q_{\zeta_{0}}=\zeta_{0}\left[\mathrm{div}\,\mathbf{v}(t.\mathbf{r})\right]^{2}$ (213)
gives
$Q_{\zeta}=\zeta(\omega)\left[\Re\!\left(\mathrm{div}\,\mathbf{v}(t.\mathbf{r})\right)\right]^{2}$ (214)
where we consider the complex velocity field with fixed frequency $\omega$
$\hat{\mathbf{v}}(t,\mathbf{r})=\mathrm{e}^{-\mathrm{i}\omega t}\,\mathbf{v}(\mathbf{r}).$ (215)
For the contribution of the temperature gradient for the loss of kinetic energy we have
$Q_{\varkappa}=\varkappa(\nabla T)^{2},$ (216)
and analogously for the ohmic dissipation in magneto-hydro-dynamics
$Q_{\omega}=\rho_{\Omega}\left(\frac{\mathrm{rot}\mathbf{B}}{\mu_{0}}\right)^{\!\!2}.$ (217)
Depending of the polarization different MHD waves have different proportion of the heating mechanism of the fluid and wave energy dissipation
$Q_{\mathrm{wave}}=Q_{\eta}+Q_{\zeta}+Q_{\varkappa}+Q_{\Omega}.$ (218)
Within this scheme we can take into account the the contribution of radiative cooling of the chromosphere plasma as negative heating
$Q_{\mathrm{rad}}=-\mathcal{P}(T)\,n_{e}n_{p},\qquad n_{e}=n_{p}=n_{\rho}\alpha^{2}.$ (219)
For solar plasma function $\mathcal{P}(T)$ was calculated and tabulated in great detail in Ref. [33]. For very low temperatures we can use the linear extrapolation in $\ln\mathcal{P}$ versus $1/T$ plot using the highest temperature $T_{0}$ ant which $\mathcal{P}_{0}=\mathcal{P}(T_{0})$ is tabulated. In such a way we obtain the approximation formula
$\mathcal{P}(T)\approx\mathcal{P}_{0}\exp(E_{1,2}(1/T_{0}-1/T)),\qquad T<T_{0},$ (220)
where
$E_{1,2}=\frac{I}{1^{2}}-\frac{I}{2^{2}}=0.75\,I$ (221)
is the difference between $2p$ and $1s$ levels for a hydrogen atom.
Knowing the energy dissipation in the short wavelength approximation we can easily calculate the wave damping rate $k^{\prime\prime}(\omega)$, see different analogous applications for calculation of wave damping [16, Eq. (25.4), Eq. (79.1)], [31, Sec. 72]. For the total heating power we have simply
$Q=Q_{\eta}+Q_{\zeta}+Q_{\varkappa}+Q_{\Omega}+Q_{\mathrm{rad}}=Q_{\mathrm{wave}}+Q_{\mathrm{rad}}.$ (222)
Now for illustration we can write the hydrodynamic equation for the fluid following the current lines [16, Eq. (5.2)]. Imagine solar chromosphere with $x$-axis in vertical direction and static solar wind velocity $U$ also in vertical direction. In this oversimplified case the energy flux is [16, Eq. (6.3)]
$q_{x}=j\tilde{w},\qquad\tilde{w}\equiv\frac{1}{2}U^{2}+w+g_{\odot}x,\qquad j\equiv\rho U,$ (223)
where $g_{\odot}$ is the gravitation acceleration. For the constant mass flux of the solar wind j(x)=j(x)=const the energy conservation gives the vertical gradient of the generalize enthalpy which includes the kinetic and potential energy of the fluid per unit mass
$\mathrm{d}_{x}\tilde{w}=\frac{Q(x)}{j}.$ (224)
For negligible dissipation $Q=0$ we obtain the Bernoulli theorem [16, Eq. (5.1)]
$\tilde{w}=\mathrm{cost},\qquad s=\mathrm{const},$ (225)
but simultaneously at these conditions the entropy per unit mass $s$ is also constant [16, Sec. 2, Sec. 5]. For the entropy of a binary cocktail
$s=\frac{S}{\rho},\qquad S=S_{e}+S_{p}+S_{0}+S_{\mathrm{He}},$ (226)
where the $S$ is the additive volume density of the entropy. According Sackur-Tetrode formula [34] and [15, Eq. (45.4)] for the partial entropy of ideal gases of electrons, protons, neutral hydrogen and helium atoms
$S_{e}$ $=\left[\ln\left(\frac{g_{e}n_{q}}{n_{e}}\right)+c_{p}\right]n_{e},\qquad g_{e}=2,\,$ (227)
$S_{p}$ $=\left[\ln\left(\frac{g_{p}n_{{}_{Q}}}{n_{e}}\right)+c_{p}\right]n_{e},\qquad g_{p}=2,\,$ (228)
$n_{{}_{Q}}\equiv\left(\frac{MT}{2\pi\hbar^{2}}\right)^{\!3/2}=n_{q}\left(\frac{M}{m}\right)^{\!3/2},\quad n_{e}=n_{p},$ (229)
$S_{0}$ $=\left[\ln\left(\frac{g_{0}n_{{}_{Q}}}{n_{0}}\right)+c_{p}\right]n_{0},\qquad g_{0}=4,\,$ (230)
$S_{\mathrm{He}}$ $=\left[\ln\left(\frac{g_{\mathrm{{}_{He}}}n_{{}_{A}}}{n_{\mathrm{{}_{He}}}}\right)+c_{p}\right]n_{\mathrm{{}_{He}}},\qquad g_{\mathrm{{}_{He}}}=1,\,$ (231)
$n_{{}_{A}}\equiv\left(\frac{AMT}{2\pi\hbar^{2}}\right)^{\!\!3/2}=n_{{}_{Q}}A^{3/2},\qquad A\approx 4,$ (232)
where $g_{e}$, $g_{p}$, $g_{0}$ and $g_{\mathrm{{}_{He}}}$ are the statistical weights of the particles, and $A$ is the mass number of the helium atom we obtain
$s=$ $\frac{1}{M^{*}}[-2\alpha\ln\alpha-(1-\alpha)\ln(1-\alpha)$ (233)
$-(1+\alpha+\overline{a}_{\mathrm{He}})\ln y-\overline{a}_{\mathrm{He}}\ln\overline{a}_{\mathrm{He}}+(1+\alpha+\overline{a}_{\mathrm{He}})c_{p}$
$+(2+3\overline{a}_{\mathrm{He}})\ln 2+c_{v}(1+\overline{a}_{\mathrm{He}})\ln(M/m)],$
where $y\equiv n_{\rho}/n_{q}$.
VI.6 A brief historical overview
A nice retrospective review by Grimus [34] on Sackur-Tetrode formula returns us to the history of the arising of quantum physics of gasses; see also the review by Flamm  [35] presented in the 20-th International Congress of History of Science, on July 25, 1997, in Liège, Belgium which we follow. Perhaps the main property of the quantum mechanics was connected in the beginning with possible discrete spectrum and quantum jumps between those levels. The probability explanation of the entropy and combinatorics calculation in the equilibrium case indispensably leads to quantization of the phase space having integer number of cells, in contemporary notations $\Delta p\Delta x/(2\pi\hbar)$. The method of quantization, i.e. introducing of discrete quantum levels of molecules in a finite volume to obtain the denumerable set of states, was used by Boltzmann in 1872 [36, L. Boltzmann: Weitere Studienüber das Wärmegleichgewicht unter Gasmolekülen, Sitzungsber. Kais. Akad. Wiss. Wien Math. Naturwiss. Classe 66 (1872) 275–370] in the paper in which he introduced the probability as a dynamic variable to derive an expression for the time derivative of entropy $\mathrm{d}_{t}S$ giving explicitly $\mathrm{d}_{t}S\geq 0$. In the present article we also use the kinetic equation Eq. (41) and because of that we return to the beginning of the story. Boltzmann actually introduced cells of finite size in phase-space spanned by the coordinates and momenta of all molecules. Much later on Born interpreted the Schrödinger wave function as an amplitude of probability $|\Psi|^{2}$. In 1877 using quantized energy states in a paper entitled “On the relation between the second law of the mechanical theory of heat and the probability calculus with respect to the theorems on thermal equilibrium” Boltzmann [37] derived the formula for the entropy which in Planck notations reads $S=k_{\mathrm{B}}\ln W.$ This formula finally led to Sackur-Tetrode formula Eq. (231) used in the present paper. Boltzmann attributed physical significance of the procedure of energy quantization. As it was recorded by Ostwald [35, 38], Boltzmann replied to a dispute with Planck at the 64-th meeting of the Deutscher Naturforscher und Ärzte, Halle, 1891 [39] with: I see no reason why energy shouldn’t also be regarded as divided atomically. Atomic, i.e. discrete in some cases. This public and recorded event was actually the insight moment (Gestalt or illumination) of the birth day of quantum mechanics (1891, Halle). In a paper in 1897 Boltzmann [40] had already suggested to Planck to use apply statistical methods for to the problem of black body radiation and oscillator equidistant levels $n=1,\,2,\,3,\,\dots$ and thermodynamic averaging $\langle n\hbar\omega\rangle=\hbar\omega/[\mathrm{e}^{\hbar\omega/T}-1$] was the first solved problem allowing comparison with an experiment on measuring the spectral density of black body radiation. The Boltzmann method of quantized cells in phase space was explicitly used by Bose in his representation of the black body radiation. The Ehrenfest method method for calculation of the Komplextion [41, 42, Eq. (28)] is given as footnote in the Landau-Lifshitz cource on theoretical physics [15, Sec. 55, Non-equilibrium Fermi and Bose gases]: $\cdot|\cdot\cdot\cdot||\cdot\cdot\cdot\cdot|\cdot\cdot$. The pictograms explain the method of calculation of 1001 different modes in which 10 identical Bose particles can be distributed in 5 Boltzmann cells in the phase $(\mathbf{p},\mathbf{x})$-space.
VI.7 Equation for hydrostatic equilibrium
At differentiation the chemical constant containing $\hbar$ disappears and we obtain in agreement with Eq. (157)
$\rho\left(\frac{\partial s}{\partial\rho}\right)_{\!\!T}$ $=-\frac{1}{M^{*}}\left[\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)\mathcal{D}\right]$ (234)
$=-\frac{1}{M^{*}}\!\left[(1+\alpha+\overline{a}_{\mathrm{He}})+\frac{(c_{v}+\iota)\,\alpha(1-\alpha)}{2-\alpha}\right]$
$=-\frac{1}{\rho}\left(\frac{\partial p}{\partial T}\right)_{\!\!\rho}=\frac{1}{\rho}\,\frac{\partial(\varepsilon-sT)}{\partial T\partial\mathcal{V}},$
which is actually the Maxwell equation for the free energy per unit mass $\varepsilon-sT$.
The vertical component of the momentum flux of the plasma fluid [16, Eq. (7.2) and Eq. (7.4)]
$\tilde{p}\equiv p+\rho\,U^{2}=p+\frac{j^{2}}{\rho}$ (235)
and momentum conservation taking into account gravitation force give
$\mathrm{d}_{x}\tilde{p}=\mathrm{d}_{x}\left[p(x)+\frac{j^{2}}{\rho(x)}\right]=-\rho(x)g_{\odot}+\mathcal{F},$ (236)
where $\mathcal{F}$ is the volume density of external for the fluid force created by wave absorption. For example, for a low frequency sound wave propagating in vertical direction we have in short wavelength approximation of homogeneous fluid
$\mathcal{F}\approx\frac{Q_{\mathrm{wave}}}{c_{0}},\qquad Q_{\mathrm{wave}}\approx 2k^{\prime\prime}(\omega)\,q_{\mathrm{wave}}.$ (237)
For negligible wind velocity and mas and wave energy flux we obtain the hydrostatic equation [16, Eq. (3.1)]
$\rho(x)=-\frac{1}{g_{\odot}}\mathrm{d}_{x}p.$ (238)
The hydrostatics of adiabatic $s(\rho,T)=\mathrm{const}$ and iso-enthalpic $w(\rho,T)=\mathrm{const}$ allows tho derive simple equations for this dissipation less motion in homogeneous gravitation field. The adiabatic condition
$\mathrm{d}_{x}s=\left(\frac{\partial s}{\partial T}\right)_{\!\!\!\rho}\mathrm{d}_{x}T+\left(\frac{\partial s}{\partial\rho}\right)_{\!\!\!T}\mathrm{d}_{x}\rho=0$ (239)
gives
$\mathrm{d}_{x}T=-\dfrac{\left(\dfrac{\partial s}{\partial\rho}\right)_{\!\!T}}{\left(\dfrac{\partial s}{\partial T}\right)_{\!\!\!\rho}}\,\mathrm{d}_{x}\rho.$ (240)
While the hydrostatic equation Eq. (238)
$\mathrm{d}_{x}p=\left(\frac{\partial p}{\partial T}\right)_{\!\!\!\rho}\mathrm{d}_{x}T+\left(\frac{\partial p}{\partial\rho}\right)_{\!\!T}\mathrm{d}_{x}\rho=-g_{\odot}\rho$ (241)
after substitution in $\mathrm{d}_{x}T$ from Eq. (240) reads
$\mathrm{d}_{x}\rho=-\dfrac{g_{\odot}\rho}{\left(\dfrac{\partial p}{\partial\rho}\right)_{\!\!\!T}-\dfrac{\left(\dfrac{\partial s}{\partial\rho}\right)_{\!\!T}}{\left(\dfrac{\partial s}{\partial T}\right)_{\!\!\!\rho}}\,\left(\dfrac{\partial p}{\partial T}\right)_{\!\!\!\rho}}.$ (242)
Hawing analytical expression for the partial derivatives Eq. (242) and Eq. (240) gives an explicit system of equations to calculate height profiles of density $\rho(x)$ end temperature.T(x)T(x). The accuracy of the numerical calculation can be evaluated whether the enthalpy Eq. (149) at this profile is constant
$w(\rho(x),T(x))\approx\mathrm{const}.$ (243)
Having reliable profiles in dissipation-less approximation we can easily evaluate the influence dissipation of energy by wave damping considered in the nest sub-section.
At known wave heating power $Q(x)$ the solution system of equations Eq. (224) and Eq. (236) gives the height profiles of the temperature $T(x)$ and density $\rho(x)$ of the chemosphere plasma.
VI.8 Hydrodynamic equations for stationary wave fluxes
In order to illustrate importance of the bulk viscosity in the physics of plasmas we will consider the simplest possible illustration of a static plasma flux flowing in a pipe with constant cross-section and heated by sound waves. This could be considered as a very rough model for heating of solar chromosphere by acoustic wave, and $U(x)$ denotes the vertical solar wind velocity, We suppose that generalized enthalpy $w(x)$ Eq. (223) and momentum flux $\Pi(x)$ Eq. (235) are static and depend only by the height $x$. In this case the energy Eq. (224) and momentum Eq. (236) conservation give the system
$\mathrm{d}_{x}\tilde{w}$ $=\left(\frac{\partial\tilde{w}}{\partial T}\right)_{\!\!\rho}\mathrm{d}_{x}T+\left(\frac{\partial\tilde{w}}{\partial\rho}\right)_{\!\!T}\mathrm{d}_{x}\rho=\frac{Q}{j},$ (244)
$\mathrm{d}_{x}\tilde{p}$ $=\left(\frac{\partial\tilde{p}}{\partial T}\right)_{\!\!\rho}\mathrm{d}_{x}T+\left(\frac{\partial\tilde{p}}{\partial\rho}\right)_{\!\!T}\mathrm{d}_{x}\rho=\mathcal{F}-\rho g_{\odot}.$ (245)
Using that
$\begin{pmatrix}\left(\dfrac{\partial\tilde{w}}{\partial T}\right)_{\!\!\rho}&\left(\dfrac{\partial\tilde{w}}{\partial\rho}\right)_{\!\!T}\\ \left(\dfrac{\partial\tilde{p}}{\partial T}\right)_{\!\!\rho}&\left(\dfrac{\partial\tilde{p}}{\partial\rho}\right)_{\!\!T}\end{pmatrix}=\begin{pmatrix}\left(\dfrac{\partial w}{\partial T}\right)_{\!\!\rho}&\qquad\left(\dfrac{\partial w}{\partial\rho}\right)_{\!\!T}-\dfrac{U^{2}}{\rho}\\ \left(\dfrac{\partial p}{\partial T}\right)_{\!\!\rho}&\qquad\left(\dfrac{\partial p}{\partial\rho}\right)_{\!\!T}-U^{2}\end{pmatrix}$, (246)
after some algebra we can express explicitly the derivatives of temperature and density
$\begin{pmatrix}\mathrm{d}_{x}T\\ \mathrm{d}_{x}\rho\end{pmatrix}=\mathsf{W}\begin{pmatrix}-g_{\odot}+Q/j\\ -g_{\odot}\rho+\mathcal{F}\end{pmatrix}$, (247)
where
$\mathsf{W}$ $\equiv\begin{pmatrix}\left(\dfrac{\partial\tilde{w}}{\partial T}\right)_{\!\!\rho}&\left(\dfrac{\partial\tilde{w}}{\partial\rho}\right)_{\!\!T}\\ \left(\dfrac{\partial\tilde{p}}{\partial T}\right)_{\!\!\rho}&\left(\dfrac{\partial\tilde{p}}{\partial\rho}\right)_{\!\!T}\end{pmatrix}^{\!\!-1}$ (248)
$=\frac{1}{\tilde{\mathcal{J}}}\begin{pmatrix}\quad\left(\dfrac{\partial p}{\partial\rho}\right)_{\!\!T}-U^{2}&\qquad-\left(\dfrac{\partial w}{\partial\rho}\right)_{\!\!T}+\dfrac{U^{2}}{\rho}\\ -\left(\dfrac{\partial p}{\partial T}\right)_{\!\!\rho}&\qquad\quad\left(\dfrac{\partial w}{\partial T}\right)_{\!\!\rho}\end{pmatrix}$.
The determinant in the denominator $\tilde{\mathcal{J}}$ is the Jacobian
$\tilde{\mathcal{J}}\equiv\frac{\partial(\tilde{w},\tilde{p})}{\partial(T,\rho)}=\left(\frac{\partial w}{\partial T}\right)_{\!\!\rho}\left(\frac{\partial\tilde{p}}{\partial\rho}\right)_{\!\!T}-\left(\frac{\partial\tilde{w}}{\partial\rho}\right)_{\!\!T}\left(\frac{\partial p}{\partial T}\right)_{\!\!\rho},$ (249)
which can be rewritten as
$\tilde{\mathcal{J}}=\mathcal{J}-U^{2}\left[\left(\frac{\partial w}{\partial T}\right)_{\!\!\mathcal{V}}-\mathcal{V}\left(\frac{\partial p}{\partial T}\right)_{\!\!\mathcal{V}}\right]=\left(c_{0}^{2}-U^{2}\right)\mathcal{C}_{v}.$ (250)
The numerical solution of these equation will be the subject of a future hydrodynamic study and in the next section we will qualitatively consider the physical conditions. The accuracy of the numerical method can be evaluated by the simple text that at $Q(x)=0$ we obtain $s(x)=\mathrm{const}.$
VII Qualitatively discussion of the applicability of the results for acoustic heating of plasmas
VII.1 Qualitative considerations
If we evaluate the shear viscosity $\eta$ of the solar chromosphere cocktail which we use as an illustration we observe huge dimensionless ratio $\mathrm{P}_{\zeta/\eta}\sim 10^{4}$ [21]. For low frequencies $f<f_{c}$ the bulk (volume) viscosity absolutely dominates and acoustic heating is created by ionization-recombination processes. Let us mention that spectral density of acoustic waves is $\propto 1/f^{\tilde{\alpha}}$ with $\tilde{\alpha}\sim 1.6-1.9.$ In the same frequency region however the absorption of Alfvén waves (AW) is negligible and they are absorbed in the transition region by the shear viscosity. The extremely small width of the transition region (TR) is perhaps determined by the positive feed-back: AW damping increases the temperature, increasing the temperature creates increasing the shear viscosity and wave damping; a process of self-induced opacity with respect to AW [43, 44]. Below the TR we suppose that the chromosphere is heated by the bulk viscosity and damping of longitudinal acoustic waves. Above the TR where these waves have already been absorbed, we have almost a Naval nozzle created by divergent magnetic force lines and this nozzle launches the supersonic solar wind.
Perhaps this scenario can be simulated by laboratory plasmas by a cocktail of alkali and noble gases, Na-Ne cocktail for example. In the initio of a pipe magneto-hydrodynamic waves can be inserted in the cold plasma. The absorption of longitudinal acoustic waves can create initial heating as in a combustion chamber. Then as in an after-burn chamber, the transversal AW are absorbed in a narrow TR by the self-induced opacity. And finally magnetic field of the enveloping solenoid can simulate a Naval nozzle for this jet engine for navigation in the solar system. However, the 3-dimensional problems and possible technical applications are far beyond the scope of our task. In the present work, we have analyzed only the bulk viscosity and sound wave damping in a homogeneous cold plasma cocktail.
VII.2 Resuming the results
Let us resume the results: We have derived the general equations for the ionic concentration oscillations created by oscillations of the density $\hat{\epsilon}_{i,a}$ Eq. (65) together with oscillation of the temperature $\hat{\epsilon}_{{}_{T}}$ Eq. (64) and electron density $\hat{\epsilon}_{e}$ Eq. (66).
These equations are significantly simplified for the cold plasmas where the temperature is much lower than the first ionization potentials $T\ll I_{a}$; see Eq. (76) and its matrix representation Eq. (86). In order to describe the acoustic heating, the propagation and absorption of sound waves with phase velocity $u_{\mathrm{s}}(\omega)$ and amplitude damping rate $k^{\prime\prime}(\omega)$, we have introduced an useful notion the generalized complex polytropic index $\hat{\gamma}(\omega)$ Eq. (97).
For an illustration we use a homogeneous plasma cocktail with parameters close to quiet solar chromosphere. In this model case we observe that the Mandelstam-Leontovich [45, 46, 23] (ML) approximation perfectly matches for the description of frequency dependent bulk viscosity. We have revealed how the ML approximation is connected with Drude [47] and Cole-Cole [22] approximations invented phenomenologically for different physical systems.
Returning to the important first ionization rate of hydrogen atom the corresponding cross-section was derived by Wannier [17] who started in 1953 from the Schroedinger equation. One can say that the bulk viscosity of cold plasmas $\zeta(\omega)$ is one of the few kinetic coefficients allowing first principle calculations. For pure hydrogen plasma the problem for the bulk viscosity $\hat{\zeta}(\omega)$ could have been solved 72 years ago.
Source code, input and output datafiles from this study are uploaded in Zenodo[48].
Acknowledgements.
The authors are thankful to Diana Bakkar, Emil Petkov, Morena Angelova for their interest in this study and to Iglika Dimitrova for the collaboration in the early stages of the research [21, 25]. The stimulating suggestions during reviewing process are highly appreciated as well. The short correspondence with Walter Grimus is also highly appreciated.